A Bibliography on Calculi for Mobile Processes

Uwe Nestmann, EPF Lausanne, Switzerland Björn Victor, Uppsala University, Sweden

May 28, 2003

Preamble

Scope. This bibliography should ideally list every publication in the context of the π-calculus [MPW92], its descendants, including higher-order forms, and implementations. In general, the listed items shall address mobility phenomena via the investigation of mathematically tractable calculi. So, this bibliography does not include just anything about process algebra (unless mobility is explicitly tackled) or mobile systems (unless a foundational calculus is involved).

Access. The main locations for accessing this bibliography are :

              http://liinwww.ira.uka.de/bibliography/Theory/pi.html
              http://www.cs.auc.dk/mobility/

The former provides a very convenient service for online searching of the information contained in the file pi.bib, the latter represents the WWW home page on Calculi for Mobile Processes and will usually provide a more up-to-date bibliography,

Updates. Any further items for inclusion and any correction of false information shall be reported to both uwe@cs.auc.dk and Bjorn.Victor@DoCS.UU.SE and will be included in future versions, if they comply with the scope of the bibliography.

TeXnics. Unpublished papers, notes, and typescripts, are currently only included if some means for anonymous electronic retrieval are known; an url-field of the respective entry in pi.bib will then point to it.

The file pi.bib makes extensive use of crossreferencing by means of crossref fields for use with bibtex. The crossref’ed entries are supplied with crossrefonly in order to enable that, like in the document pi-bib.dvi (this document), corresref’ed Proceedings etc do not have to be listed themselves.

Pretty-printing of pi.bib and the generation of its citekeys is done by using bibtool-2.44. Several abbreviations are used by default in the current bibliography (mainly in order to reduce

the size of the dvi-file), like e.g. writing LNCS instead of Lecture Notes in Computer Science. The alternative extensions are contained in pi.bib as @COMMENT{…}.

   For nicer TeXing of url’s, we use the package ftp://ftp.tex.ac.uk/tex-archive/macros/

latex/contrib/other/misc/url.sty, which intelligently performs hyphenation of url’s.

References

[AA92] R. M. Amadio and O. Aït-Mohamed. An Analysis of π-Calculus Bisimulation. Technical Report ECRC-94-2, European Computer-Industry Research Centre, Munich, Jan. 1992.

[AB01] M. Abadi and B. Blanchet. Secrecy Types for Asymmetric Communication. In F. Honsell and M. Miculan, eds, Proceedings of FoSSaCS 2001, volume 2030 of LNCS, pages 25–41. Springer, 2001.

[Aba99] M. Abadi. Secrecy by Typing in Security Protocols. Journal of the ACM, 46(5):749– 786, Sept. 1999. An abstract appeared in the Proceedings of TACS ’97, volume 1281 of LNCS.

[ABL99] R. M. Amadio, G. Boudol and C. Lhoussaine. The Receptive Distributed Pi-Calculus (Extended Abstract). In C. Pandu Rangan, V. Raman and R. Ramanujam, eds, Proceedings of FSTTCS ’99, volume 1738 of LNCS, pages 304–315. Springer, Dec. 1999.

[Abr94] S. Abramsky. Proofs as Processes. Theoretical Computer Science, 135:5–9, 1994. An Introduction to [BS94a].

[ACS98] R. M. Amadio, I. Castellani and D. Sangiorgi. On Bisimulations for the Asynchronous π-Calculus. Theoretical Computer Science, 195(2):291–324, 1998. An extended abstract appeared in Proceedings of CONCUR ’96, LNCS 1119: 147–162.

[AD95] R. M. Amadio and M. Dam. Reasoning about Higher-Order Processes. In P. D. Mosses, M. Nielsen and M. I. Schwarzbach, eds, Proceedings of TAPSOFT ’95, volume 915 of LNCS, pages 202–216. Springer, 1995. Full version as SICS Research Report RR:94/18, October 1994.

[AD96] R. M. Amadio and M. Dam. Toward a Modal Theory of Types for the π-Calculus. In B. Jonsson, ed, Proceedings of FTRTFT’96, volume 1135 of LNCS, pages 347–365. Springer, 1996.

[AF01] M. Abadi and C. Fournet. Mobile Values, New Names, and Secure Communication. In Proceedings of POPL ’01, pages 104–115. ACM, Jan. 2001.

[AFG98] M. Abadi, C. Fournet and G. Gonthier. Secure Implementation of Channel Abstractions. In Proceedings of LICS ’98. IEEE, Computer Society Press, July 1998.

[AFG99] M. Abadi, C. Fournet and G. Gonthier. A Top-Down Look at a Secure Message. In C. Pandu Rangan, V. Raman and R. Ramanujam, eds, Proceedings of FSTTCS ’99, volume 1738 of LNCS, pages 122–141. Springer, Dec. 1999.

[AFG00] M. Abadi, C. Fournet and G. Gonthier. Authentication primitives and their compilation. In Proceedings of POPL ’00, pages 302–315. ACM, Jan. 2000.

[AG97] M. Abadi and A. D. Gordon. Reasoning about Cryptographic Protocols in the Spi Calculus. In A. Mazurkiewicz and J. Winkowski, eds, Proceedings of CONCUR ’97, volume 1243 of LNCS, pages 59–73. Springer, 1997.

[AG98] M. Abadi and A. D. Gordon. A Bisimulation Method for Cryptographic Protocols. Nordic Journal of Computing, 5(4):267–303, Winter 1998. An extended abstract appeared in the Proceedings of ESOP ’98, LNCS 1381, pages 12–26.

[AG99] M. Abadi and A. D. Gordon. A Calculus for Cryptographic Protocols: The Spi Calculus. Journal of Information and Computation, 148(1):1–70, 1999. An extended abstract appeared in the Proceedings of the Fourth ACM Conference on Computer and Communications Security (Zürich, April 1997). An extended version of this paper appears as Research Report 149, Digital Equipment Corporation Systems Research Center, January 1998, and, in preliminary form, as Technical Report 414, University of Cambridge Computer Laboratory, January 1997.

[Aït95] O. Aït-Mohamed. Mechanizing a pi-calculus equivalence in HOL. In E. T. Schubert, P. J. Windley and J. Alves Foss, eds, Proceedings of HOL ’95, volume 971 of LNCS, pages 1–16. Springer, 1995. See also Rapport de Recherche RR-2412, INRIA-Lorraine, 1994.

[Aït96] O. Aït-Mohamed. The theory of the pi-calculus in HOL. PhD thesis, Henri Poincaré University Nancy I, France, July 1996.

[AL00] R. M. Amadio and D. Lugiez. On the Reachability Problem in Cryptographic Protocols. In C. Palamidessi, ed, Proceedings of CONCUR 2000, volume 1877 of LNCS, pages 380–394. Springer, Aug. 2000.

[ALT95] R. M. Amadio, L. Leth and B. Thomsen. From a Concurrent λ-Calculus to the π-Calculus. In H. Reichel, ed, Proceedings of FCT ’95, volume 965 of LNCS, pages 106–115. Springer, 1995. Full version as Technical Report ECRC-95-18.

[Ama92] R. M. Amadio. A uniform presentation of CHOCS and π-calculus. Rapport de Recherche 1726, INRIA-Lorraine, Nancy, 1992.

[Ama93] R. M. Amadio. On the reduction of CHOCS bisimulation to π-calculus bisimulation. In E. Best, ed, Proceedings of CONCUR ’93, volume 715 of LNCS, pages 112–126. Springer, 1993. Extended version as Rapport de Recherche, INRIA-Lorraine, 1993.

[Ama94] R. M. Amadio. Translating Core Facile. Technical Report ECRC-94-3, European Computer-Industry Research Centre, München, Feb. 1994.

[Ama97] R. M. Amadio. An Asynchronous Model of Locality, Failure, and Process Mobility. In D. Garlan and D. Le Metayer, eds, Proceedings of COORDINATION ’97, volume 1282 of LNCS. Springer, 1997. Extended version as Rapport de Recherche RR-3109, INRIA Sophia-Antipolis, 1997.

[Ama00] R. M. Amadio. On modelling mobility. Theoretical Computer Science, 240(1):147–176, 2000.

[Amb91] S. J. Ambler. A de Bruijn notation for the π-calculus. Technical Report 569, Dept. of Computer Science, Queen Mary and Westfield College, London, May 1991.

[AP94] R. M. Amadio and S. Prasad. Localities and Failures. In P. S. Thiagarajan, ed, Proceedings of FSTTCS ’94, volume 880 of LNCS, pages 205–216. Springer, 1994. Full version as Technical Report 94-18, ECRC Munich, 1994.

[AP98] R. M. Amadio and S. Prasad. Modelling IP Mobility. In D. Sangiorgi and R. de Simone, eds, Proceedings of CONCUR ’98, volume 1466 of LNCS, pages 301–316. Springer, Sept. 1998.

[AP99] R. M. Amadio and S. Prasad. The game of the Name in Cryptographic Tables. In P. S. Thiagarajan and R. Yap, eds, Proceedings of ASIAN ’99, volume 1742 of LNCS, pages 15–26. Springer, Dec. 1999.

[AZ84] E. Astesiano and E. Zucca. Parametric Channels via Label Expressions in CCS. Theoretical Computer Science, 33:45–64, 1984.

[BB92] G. Berry and G. Boudol. The chemical abstract machine. Theoretical Computer Science, 96:217–248, 1992.

[BB02a] M. Boreale and M. Buscemi. A Framework for the Analysis of Security Protocols. In Proceedings of CONCUR’02, 2002. To appear.

[BB02b] M. Boreale and M. G. Buscemi. A Framework for the Analysis of Security Protocols. In L. Brim, P. Jančar, M. Křetinský and A. Kučera, eds, Proceedings of CONCUR 2002, volume 2421 of LNCS, pages 483–498. Springer, Aug. 2002.

[BBP95] R. Banach, J. Balász and G. Papadoupoulos. A Translation of the Pi-Calculus into MONSTR. Journal of Universal Computer Science, 1(6):339–398, 1995.

[BC01] M. Bugliesi and G. Castagna. Secure Safe Ambients. In Proceedings of POPL ’01, pages 222–235. ACM, Jan. 2001.

[BCC00] M. Bugliesi, G. Castagna and S. Crafa. Typed Mobile Objects. In C. Palamidessi, ed, Proceedings of CONCUR 2000, volume 1877 of LNCS, pages 504–520. Springer, Aug. 2000.

[BCDP95] A. Bianchi, S. Coluccini, P. Degano and C. Priami. An Efficient Verifier of Truly Concurrent Properties. In V. Malyshkin, ed, Proceedings of PaCT ’95, volume 964 of LNCS, pages 36–50. Springer, 1995.

[BD95] M. Boreale and R. De Nicola. Testing Equivalences for Mobile Processes. Journal of Information and Computation, 120:279–303, 1995. Available as Report SI 92 RR 04, Università “La Sapienza” di Roma; an extended abstract appeared in Proceedings of CONCUR ’92, LNCS 630.

[BD96] M. Boreale and R. De Nicola. A Symbolic Semantics for the π-Calculus. Journal of Information and Computation, 126(1):34–52, 1996. Available as Report SI 94 RR 04, Università “La Sapienza” di Roma; an extended abstract appeared in Proceedings of CONCUR ’94, pages 299–314, LNCS 836.

[BD99] G. Boudol and S. Dal-Zilio. An Interpretation of Extensible Objects. In G. Ciobanu and G. Păun, eds, Proceedings of FCT ’99: Fundamentals of Computation Theory, 12th International Symposium (Iasi, Romania, August 30 - September 3, 1999), volume 1684 of LNCS. Springer, 1999.

[BDFP02] C. Bodei, P. Degano, R. Focardi and C. Priami. Primitives for Authentication in Process Algebras. Theoretical Computer Science, 283(2):271–304, 2002.

[BDNN98] C. Bodei, P. Degano, F. Nielson and H. R. Nielson. Control Flow Analysis for the Pi-Calculus. In D. Sangiorgi and R. de Simone, eds, Proceedings of CONCUR ’98, volume 1466 of LNCS, pages 84–98. Springer, Sept. 1998.

[BDNN99] C. Bodei, P. Degano, F. Nielson and H. R. Nielson. Static Analysis of Processes for No Read-Up and No Write-Down. In W. Thomas, ed, Proceedings of FoSSaCS ’99, volume 1578 of LNCS, pages 120–134. Springer, 1999.

[BDP96a] C. Bodei, P. Degano and C. Priami. Mobile Processes With a Distributed Envirionent. In F. Meyer auf der Heide and B. Monien, eds, Proceedings of ICALP ’96, volume 1099 of LNCS, pages 490–501. Springer, 1996. Full version as Università di Pisa Technical Report “Handling Locally Names of Mobile Agents”, 1996.

[BDP^{+}96b] R. Borgia, P. Degano, C. Priami, L. Leth and B. Thomsen. Understanding Mobile Agents via a Non Interleaving Semantics for Facile. In R. Cousot and D. A. Schmidt, eds, Proceedings of SAS ’96, volume 1145 of LNCS, pages 98–112. Springer, 1996. Extended version as Technical Report ECRC-96-4, 1996.

[BDP98] M. Boreale, R. De Nicola and R. Pugliese. Asynchronous Observations of Processes. In M. Nivat, ed, Proceedings of FoSSaCS ’98, volume 1378 of LNCS, pages 95–109. Springer, 1998.

[BDP99a] M. Boreale, R. De Nicola and R. Pugliese. Proof Techniques for Cryptographic Processes. In Proceedings of LICS ’99, pages 157–166. IEEE, Computer Society Press, July 1999.

[BDP99b] M. Boreale, R. De Nicola and R. Pugliese. A Theory of “May” Testing for Asynchronous Languages. In W. Thomas, ed, Proceedings of FoSSaCS ’99, volume 1578 of LNCS, pages 165–179. Springer, 1999.

[BDP00a] M. Boreale, R. De Nicola and R. Pugliese. Process Algebraic Analysis of Cryptographic Protocols. In Proceedings of FORTE/PSTV’00: Formal Description Techniques (FORTE XIII) & Protocol Specification, Testing, and Verification (PSTV XX), volume 183 of IFIP Conference Proceedings, 2000.

[BDP00b] L. Brodo, P. Degano and C. Priami. A Tool for Quantitative Analysis of pi-Calculus Processes. In R. Gorrieri, ed, Proceedings of PAPM’00, Geneva, 2000. Carleton Scientific.

[BDP01] C. Bodei, P. Degano and C. Priami. Names of the pi-Calculus Agents Handled Locally. Theoretical Computer Science, 253(2):155–184, Feb. 2001.

[BDP02] M. Boreale, R. De Nicola and R. Pugliese. Proof Techniques for Cryptographic Processes. SIAM Journal on Computing, 31(3):947–986, 2002.

[BFL98] M. Boreale, C. Fournet and C. Laneve. Bisimulations for the Join-Calculus. In D. Gries and W.-P. de Roever, eds, Proceedings of PROCOMET ’98, pages 68–86, 1998. Full version to appear in TCS.

[BG95] N. Busi and R. Gorrieri. A Petri Net Semantics for π-Calculus. In I. Lee and S. A. Smolka, eds, Proceedings of CONCUR ’95, volume 962 of LNCS, pages 145–159. Springer, 1995.

[BG02a] M. Boreale and D. Gorla. On Compositional Reasoning in the Spi-Calculus. In M. Nielsen and U. H. Engberg, eds, Proceedings of FoSSaCS 2002, volume 2303 of LNCS, pages 67–81. Springer, Apr. 2002.

[BG02b] M. Boreale and D. Gorla. Process Calculi and the Verification of Security Properties. Journal of Telecommunication and Information Technology — Special Issue on Cryptographic Protocol Verification, (4/02), 2002.

[BH00] M. Berger and K. Honda. The Two-Phase Commitment Protocol in an Extended pi-Calculus. In L. Aceto and B. Victor, eds, Proceedings of EXPRESS ’00, volume 39.1 of ENTCS. Elsevier Science Publishers, 2000.

[BHY01] M. Berger, K. Honda and N. Yoshida. Sequentiality and the pi-Calculus. In Proceedings of TCLA 2001, volume 2044 of LNCS, pages 29–45. Springer, 2001.

[BL00] G. Boudol and C. Laneve. λ-Calculus, Multiplicites and the π-Calculus. In G. Plotkin, C. Stirling and M. Tofte, eds, Proof, Language and Interaction: Essays in Honour of Robin Milner, Foundations of Computing, May 2000.

[BLM02] R. Bruni, C. Laneve and U. Montanari. Orchestrating Transactions in Join Calculus. In L. Brim, P. Jančar, M. Křetinský and A. Kučera, eds, Proceedings of CONCUR 2002, volume 2421 of LNCS, pages 321–337. Springer, Aug. 2002.

[BM02] M. G. Buscemi and U. Montanari. A First Order Coalgebraic Model of π-Calculus Early Observational Equivalence. In L. Brim, P. Jančar, M. Křetinský and A. Kučera, eds, Proceedings of CONCUR 2002, volume 2421 of LNCS, pages 449–465. Springer, Aug. 2002.

[BN02] J. Borgström and U. Nestmann. On Bisimulations for the Spi Calculus. In H. Kirchner and C. Ringeissen, eds, Proceedings of AMAST 2002, volume 2422 of LNCS, pages 287–303. Springer, 2002.

[BO95] S. Brock and G. Ostheimer. Process Semantics of Graph Reduction. In I. Lee and S. A. Smolka, eds, Proceedings of CONCUR ’95, volume 962 of LNCS, pages 471–485. Springer, 1995.

[Bod00] C. Bodei. Security Issues in Process Calculi. PhD thesis, Università degli Studi di Pisa, Mar. 2000. Available as TD-2/00.

[Bor95] M. Boreale. Process Algebraic Theories for Mobile Systems. PhD thesis, Università degli Studi di Roma “La Sapienza”, 1995. Available as PhD VII-95-2.

[Bor98] M. Boreale. On the Expressiveness of Internal Mobility in Name-Passing Calculi. Theoretical Computer Science, 195(2):205–226, 1998. An extended abstract appeared in Proceedings of CONCUR ’96, LNCS 1119: 163–178.

[Bor01] M. Boreale. Symbolic Trace Analysis of Cryptographic Protocols. In Proceedings of ICALP 2001, volume 2076 of LNCS, pages 667–681. Springer, July 2001.

[Bou89] G. Boudol. Towards a Lambda-Calculus for Concurrent and Communicating Systems. In J. Díaz and F. Orejas, eds, Proceedings of TAPSOFT ’89, Volume 1, volume 351 of LNCS, pages 149–161. Springer, 1989.

[Bou92] G. Boudol. Asynchrony and the π-calculus (Note). Rapport de Recherche 1702, INRIA Sophia-Antipolis, May 1992.

[Bou94] G. Boudol. Some chemical abstract machines. In J. W. de Bakker, W.-P. de Roever and G. Rozenberg, eds, A Decade of Concurrency 1993 (REX Workshop), volume 803 of LNCS, pages 92–123. Springer, 1994.

[Bou97] G. Boudol. Typing the Use of Resources in a Concurrent Calculus. In R. K. Shyamasundar and K. Ueda, eds, Proceedings of ASIAN ’97, volume 1345 of LNCS, pages 239–253. Springer, 1997.

[Bou98] G. Boudol. The π-Calculus in Direct Style. Higher-Order and Symbolic Computation, 11:177–208, 1998. Previously published in the Proceedings of POPL ’97, pages 228– 241.

[BS94a] G. Bellin and P. Scott. On the π-Calculus and Linear Logic. Theoretical Computer Science, 135:11–65, 1994. Also published as LFCS report ECS-LFCS-92-232, LFCS, University of Edinburgh.

[BS94b] J. Bergstra and G. Stefanescu. Network Algebra for Synchronous and Asynchronous Dataflow. Logic Group Preprint Series 122, Department of Philosophy, Utrecht University, The Netherlands, Oct. 1994.

[BS98a] M. Boreale and D. Sangiorgi. Bisimulation in name-passing calculi without matching. In Proceedings of LICS ’98. IEEE, Computer Society Press, July 1998.

[BS98b] M. Boreale and D. Sangiorgi. A Fully Abstract Semantics for Causality in the pi-Calculus. Acta Informatica, 35(5):353–400, 1998. An extended abstract appeared in the Proceedings of STACS ’95, volume 900 of LNCS. A long version appeared as report ECS–LFCS–94–297, University of Edinburgh.

[BS98c] M. Boreale and D. Sangiorgi. Some Congruence Properties for π-calculus Bisimilarities. Theoretical Computer Science, 198(1–2):159–176, 1998. Previous version as Rapport de Recherche RR-2870, INRIA Sophia-Antipolis, 1996.

[BS01] M. G. Buscemi and V. Sassone. High-Level Petri Nets as Type Theories in the Join Calculus. In F. Honsell and M. Miculan, eds, Proceedings of FoSSaCS 2001, volume 2030 of LNCS, pages 104–120. Springer, 2001.

[Cai96] L. Caires. A language for the logical specification of processes and relations. In M. Hanus, ed, Proceedings of the Algebraic and Logic Programming International Conference ALP’96, number 1139 in Lecture Notes in Computer Science, pages 150– 164. Springer-Verlag, 1996.

[Cai99] L. Caires. A Model for Declarative Programming and Specification with Concurrency and Mobility. PhD thesis, Universidade Nova de Lisboa, 1999.

[CC02] L. Caires and L. Cardelli. A Spatial Logic for Concurrency (Part II). In L. Brim, P. Jančar, M. Křetinský and A. Kučera, eds, Proceedings of CONCUR 2002, volume 2421 of LNCS, pages 209–225. Springer, Aug. 2002.

[CDG^{+}01] W. Charatonik, S. Dal Zilio, A. D. Gordon, S. Mukhopadhyay, and J.-M. Talbot. The Complexity of Model Checking Mobile Ambients. In F. Honsell and M. Miculan, eds, Proceedings of FoSSaCS 2001, volume 2030 of LNCS, pages 52–167. Springer, 2001.

[CG99] L. Cardelli and A. D. Gordon. Types for Mobile Ambients. In Proceedings of POPL ’99, pages 79–92. ACM, 1999.

[CG00a] L. Cardelli and A. D. Gordon. Anytime, Anywhere: Modal Logics for Mobile Ambients. In Proceedings of POPL ’00, pages 365–377. ACM, Jan. 2000.

[CG00b] L. Cardelli and A. D. Gordon. Mobile Ambients. Theoretical Computer Science, 240(1):177–213, 2000. An extended abstract appeared in Proceedings of FoSSaCS ’98 : 140–155.

[CG01a] L. Cardelli and G. Ghelli. A Query Language Based on the Ambient Logic. In D. Sands, ed, Proceedings of ESOP 2001, volume 2028 of LNCS, pages 1–22. Springer, 2001.

[CG01b] L. Cardelli and A. D. Gordon. Logical Properties of Name Restriction. In Proceedings of TCLA 2001, volume 2044 of LNCS, pages 46–60. Springer, 2001.

[CGG99] L. Cardelli, G. Ghelli and A. D. Gordon. Mobility Types for Mobile Ambients. In J. Wiederman, P. van Emde Boas and M. Nielsen, eds, Proceedings of ICALP ’99, volume 1644 of LNCS, pages 230–239. Springer, July 1999.

[CGG00a] L. Cardelli, G. Ghelli and A. D. Gordon. Ambient Groups and Mobility Types. In J. van Leeuwen, O. Watanabe, M. Hagiya, P. Mosses and T. Ito, eds, Proceedings of TCS 2000, volume 1872 of LNCS, pages 333–347. IFIP, Springer, Aug. 2000.

[CGG00b] L. Cardelli, G. Ghelli and A. D. Gordon. Secrecy and Group Creation. In C. Palamidessi, ed, Proceedings of CONCUR 2000, volume 1877 of LNCS, pages 365–379. Springer, Aug. 2000.

[CGG02] L. Cardelli, G. Ghelli and A. D. Gordon. Types for the Ambient Calculus. Submitted for publication, 2002.

[CH98] I. Castellani and M. Hennessy. Testing Theories for Asynchronous Languages. In V. Arvind and R. Ramanujam, eds, Proceedings of FSTTCS ’98, volume 1530 of LNCS, pages 90–101. Springer, Dec. 1998.

[CM98] L. Caires and L. Monteiro. Verifiable and Executable Specifications of Concurrent Objects in L_{π}. In C. Hankin, ed, Proceedings of ESOP ’98, volume 1381 of LNCS, pages 42–56. Springer, 1998. To appear in TCS.

[Cor02] V. Cortier. Observational Equivalence and Trace Equivalence in an Extension of Spicalculus. Application to Cryptographic Protocol Analysis. Technical Report LSV-02-3, École Nationale Supérieure de Cachan, France, 2002.

[CR98] G. Ciobanu and M. Rotaru. Faithful π-nets. A graphical representation of the asynchronous π-calculus. In P. Jancar and M. Kretinsky, eds, Proceedings of MFCS ’98 Workshop on Concurrency, volume 18 of ENTCS. Elsevier Science Publishers, 1998.

[Cra92] C. Crasemann. πλ-Kalküle für Prozesse und Funktionen. Dissertation, Christian-Albrechts-Universität Kiel, Feb. 1992. Insitut für Informatik und Praktische Mathematik, Bericht Nr. 9302 (in German).

[CS01] T. Chothia and I. Stark. A Distributed Calculus with Local Areas of Communication. In HLCL ’00: Proceedings of the 4th International Workshop on High-Level Concurrent Languages, volume 41.2 of Electronic Notes in Theoretical Computer Science. Elsevier, 2001.

[CSW97] G. L. Cattani, I. Stark and G. Winskel. Presheaf Models for the Pi-Calculus. In E. Moggi and G. Rosolini, eds, Proceedings of CTCS ’97, volume 1290 of LNCS, pages 106–126. Springer, 1997.

[Dal94] S. Dal-Zilio. Modèles du π-calcul. Structures d’Actions et Relations avec l’Équivalence Par Permutations. Mémoire de D.E.A., École Normale Superieure de Lyon, INRIA Sophia-Antipolis, 1994. (in French).

[Dal97] S. Dal-Zilio. Implicit Polymorphic Type System for the Blue Calculus. Rapport de Recherche RR-3244, INRIA Sophia-Antipolis, Sept. 1997.

[Dal98] S. Dal-Zilio. Quiet and Bouncing Objects: Two Migration Abstractions in a Simple Distributed Blue Calculus. In H. Hüttel and U. Nestmann, eds, Proceedings of SOAP ’98, volume NS-98-5 of BRICS Notes Series, pages 35–42, 1998.

[Dal99] S. Dal-Zilio. Calcul bleu: types et objets. Thèse d’état, Université de Nice - Sophia-Antipolis, 1999.

[Dal00a] S. Dal-Zilio. An Interpretation of Typed Concurrent Objects in the Blue Calculus. In J. van Leeuwen, O. Watanabe, M. Hagiya, P. Mosses and T. Ito, eds, Proceedings of TCS 2000, volume 1872 of LNCS, pages 409–424. IFIP, Springer, Aug. 2000.

[Dal00b] S. Dal-Zilio. Spatial Congruence for Ambients Is Decidable. In J. He and M. Sato, eds, Proceedings of ASIAN 2000, volume 1961 of LNCS, pages 88–103. Springer, Dec. 2000.

[Dam96] M. Dam. Model Checking Mobile Processes. Journal of Information and Computation, 129(1):35–51, 1996. Also available as Research Report R94:01, SICS, Sweden. A summary appeared in Proceedings of CONCUR ’93 , LNCS 715.

[Dam97] M. Dam. On the Decidability of Process Equivalences for the π-Calculus. Theoretical Computer Science, 183:215–228, 1997. Previously published as SICS Research Report R94:20, 1994.

[Dam98a] M. Dam. Proving properties of dynamic process networks. Journal of Information and Computation, 140(2):95–114, Feb. 1998.

[Dam98b] M. Dam. Proving Trust in Systems of Second-Order Processes (Extended Abstract). In Proceedings of HICSS 31, volume VII, pages 255–264. IEEE Computer Science, 1998.

[Dam01] M. Dam. Proof Systems for pi-Calculus Logics. In R. de Queiroz, ed, Logic for Concurrency and Synchronisation, Studies in Logic and Computation. Oxford University Press, 2001. To appear.

[DDG99] F. Damiani, M. Dezani-Ciancaglini and P. Giannini. A Filter Model for Mobile Processes. Mathematical Structures in Computer Science, 9(1):63–101, 1999.

[Des00] J. Despeyroux. A Higher-Order Specification of the π-Calculus. In J. van Leeuwen, O. Watanabe, M. Hagiya, P. Mosses and T. Ito, eds, Proceedings of TCS 2000, volume 1872 of LNCS, pages 425–439. IFIP, Springer, Aug. 2000.

[DFPV00] R. De Nicola, G. Ferrari, R. Pugliese and B. Veneri. Types for Access Control. Theoretical Computer Science, 240(1):215–254, 2000.

[DG00] S. Dal-Zilio and A. D. Gordon. Region Analysis and a pi-Calculus with Groups. In M. Nielsen and B. Rovan, eds, Proceedings of MFCS 2000, volume 1893 of LNCS, pages 409–424. Springer, 2000.

[DLB00] P. Degano, F. Levi and C. Bodei. Safe Ambients: Control Flow Analysis and Security. In J. He and M. Sato, eds, Proceedings of ASIAN 2000, volume 1961 of LNCS, pages 199–214. Springer, Dec. 2000.

[DLP97] P. Degano, J.-V. Loddo and C. Priami. Mobile Processes with Local Clocks. In M. Dam, ed, Proceedings of the Fifth LOMAPS Workshop, volume 1192 of LNCS. Springer, 1997.

[DP95] P. Degano and C. Priami. Causality for Mobile Processes. In Z. Fülöp and F. Gécseg, eds, Proceedings of ICALP ’95, volume 944 of LNCS, pages 660–671. Springer, 1995.

[DP96] P. Degano and C. Priami. Enhanced Operational Semantics. ACM Computing Surveys, 28(2):352–354, 1996.

[DP97] P. Degano and C. Priami. Non Interleaving Semantics for Mobile Processes. Theoretical Computer Science, 1997. To appear.

[DPLT97] P. Degano, C. Priami, L. Leth and B. Thomsen. Analysis of Facile programs: a case study. In M. Dam, ed, Proceedings of the Fifth LOMAPS Workshop, volume 1192 of LNCS. Springer, 1997.

[DPLT99] P. Degano, C. Priami, L. Leth and B. Thomsen. Causality for Debugging Mobile Agent. Acta Informatica, 36(5):335–374, 1999.

[DSV00] L. Durante, R. Sisto and A. Valenzano. A State-Exploration Technique for Spi-Calculus Testing-Equivalence Verification. In Proceedings of FORTE/PSTV 2000 (Pisa, Italy), volume 183 of IFIP Conference Proceedings, pages 155–170, Oct. 2000.

[EG99] J. Engelfriet and T. Gelsema. Multisets and Structural Congruence of the π-calculus with Replication. Theoretical Computer Science, 211(1–2):311–337, 1999. Previously published as Leiden University Report 95-02, 1995.

[EG01] J. Engelfriet and T. Gelsema. Structural Inclusion in the pi-Calculus with Replication. Theoretical Computer Science, 258(1–2):131–168, 2001.

[EHHO99] A. S. Elkjær, M. Höhle, H. Hüttel and K. Overgård. Towards Automatic Bisimilarity Checking in the Spi Calculus. In C. S. Calude and M. J. Dinneen, eds, Combinatorics, Computation & Logic, volume 21(3) of Australian Computer Science Communications, pages 175–189. Springer-Verlag Singapore Pte. Ltd., Jan. 1999. As part of the Australian Computer Science Week, January 18–21, 1999.

[EN86] U. Engberg and M. Nielsen. A Calculus of Communicating Systems with Labelpassing. Technical Report DAIMI PB-208, Comp. Sc. Department, Univ. of Aarhus, Denmark, 1986.

[Eng96] J. Engelfriet. A Multiset Semantics for the π-calculus with Replication. Theoretical Computer Science, 153(1–2):65–94, 1996. An extract appeared in Proceedings of CONCUR ’93, LNCS 715, and as Leiden University Report 94-26, August 1994.

[EO95] L.-H. Eriksson and F. Orava. Analysing a π-calculus Specification of a Cache-Coherence Protocol. Technical Report SICS RR95-04, Swedish Institute of Computer Science, 1995.

[EP93] S. Eisenbach and R. Patterson. π-Calculus Semantics for the Concurrent Configuration Language Darwin. In Proceedings of the 26th Annual Hawaii International Conference on System Sciences, volume 2. IEEE Computer Society Press, 1993.

[FA] M. Fiore and M. Abadi. Computing Symbolic Models for Verifying Cryptographic Protocols.

[Fer00a] J. Feret. Confidentiality Analysis of Mobile Systems. In J. Palsberg, ed, Proceedings of SAS 2000, volume 1824 of LNCS, pages 135–154. Springer, 2000.

[Fer00b] J. Feret. Occurrence Counting Analysis for the pi-Calculus. In P. Cousot, E. Goubault, J. Gunawardena, M. Herlihy, M. Raussen and V. Sassone, eds, Electronic Notes in Theoretical Computer Science, volume 39.2 of ENTCS, 2000.

[Fer01] J. Feret. Abstract Interpretation-Based Static Analysis of Mobile Ambients. In Proceedings of SAS ’01, LNCS. Springer, 2001. To appear.

[FFG^{+}97] G. Ferrari, G. Ferro, S. Gnesi, U. Montanari, M. Pistore and G. Ristori. An Automata Based Verification Environment for Mobile Processes. In E. Brinksma, ed, Proceedings of TACAS ’97, volume 1217 of LNCS, pages 275–289. Springer, Apr. 1997.

[FG96] C. Fournet and G. Gonthier. The Reflexive Chemical Abstract Machine and the Join-Calculus. In Proceedings of POPL ’96, pages 372–385. ACM, Jan. 1996.

[FG98] C. Fournet and G. Gonthier. A Hierarchy of Equivalences for Asynchronous Calculi. In K. G. Larsen, S. Skyum and G. Winskel, eds, Proceedings of ICALP ’98, volume 1443 of LNCS, pages 844–855. Springer, July 1998.

[FGL^{+}96] C. Fournet, G. Gonthier, J.-J. Lévy, L. Maranget and D. Rémy. A Calculus of Mobile Agents. In U. Montanari and V. Sassone, eds, Proceedings of CONCUR ’96, volume 1119 of LNCS, pages 406–421. Springer, 1996.

[FGM^{+}98] G. Ferrari, S. Gnesi, U. Montanari, M. Pistore and G. Ristori. Verifying Mobile Processes in the HAL Environment. In A. J. Hu and M. Y. Vardi, eds, Proceedings of CAV ’98, volume 1427 of LNCS. Springer, 1998. Tool Poster.

[FHJ02] U. Frendrup, H. Hüttel and J. N. Jensen. Modal Logics for Cryptographic Processes. In U. Nestmann and P. Panangaden, eds, Proceedings of EXPRESS 2002, volume 68 of ENTCS. Elsevier Science Publishers, 2002.

[FLMR97] C. Fournet, C. Laneve, L. Maranget and D. Rémy. Implicit Typing à la ML for the join-calculus. In A. Mazurkiewicz and J. Winkowski, eds, Proceedings of CONCUR ’97, volume 1243 of LNCS, pages 196–212. Springer, 1997.

[FLMR00] C. Fournet, C. Laneve, L. Maranget and D. Rémy. Inheritance in the Join Calculus. In S. Kapoor and S. Prasad, eds, Proceedings of FSTTCS 2000, volume 1974 of LNCS, pages 397–408. Springer, Dec. 2000.

[FLS00] C. Fournet, J.-J. Lévy and A. Schmitt. An Asynchronous Distributed Implementation fo Mobile Ambients. In J. van Leeuwen, O. Watanabe, M. Hagiya, P. Mosses and T. Ito, eds, Proceedings of TCS 2000, volume 1872 of LNCS, pages 348–364. IFIP, Springer, Aug. 2000.

[FM94] G. Ferrari and U. Montanari. A Tile-Based Coordination View of Asynchronous πcalculus. In I. Prívara and P. Ružička, eds, Proceedings of MFCS ’97, volume 1295 of LNCS. Springer, Aug. 1994.

[FMP02] G. Ferrari, U. Montanari and M. Pistore. Minimizing Transition Systems for Name Passing Calculi: A Co-algebraic Formulation. In M. Nielsen and U. H. Engberg, eds, Proceedings of FoSSaCS 2002, volume 2303 of LNCS, pages 129–143. Springer, Apr. 2002.

[FMQ94] G. Ferrari, U. Montanari and P. Quaglia. A π-calculus with explicit substitutions: the late Semantics. In I. Prívara, B. Rovan and P. Ružička, eds, Proceedings of MFCS ’94, volume 841 of LNCS, pages 342–351. Springer, Aug. 1994. Full version as Università di Pisa Report, 1994.

[FMQ95a] G. Ferrari, G. Modoni and P. Quaglia. Towards a Semantic-Based Environment for the π-Calculus. In Proceedings of 5th Italian Conference on Theoretical Computer Science (ICTCS-95). World Scientific, 1995.

[FMQ95b] G. Ferrari, U. Montanari and P. Quaglia. The Weak Late π-Calculus Semantics as Observation Equivalence. In I. Lee and S. A. Smolka, eds, Proceedings of CONCUR ’95, volume 962 of LNCS, pages 57–71. Springer, 1995.

[FMQ96] G. Ferrari, U. Montanari and P. Quaglia. A π-calculus with explicit substitutions. tcs, 168(1):53–103, 1996.

[FMS96] M. Fiore, E. Moggi and D. Sangiorgi. A Fully-Abstract Model for the π-calculus (Extended Abstract). In Proceedings of LICS ’96, pages 43–54. IEEE, Computer Society Press, July 1996.

[Fou98] C. Fournet. The Join-Calculus: A Calculus for Distributed Mobile Programming. PhD thesis, École Polytechnique, Paris, France, 1998.

[Fu97] Y. Fu. A proof-theoretical approach to communication. In P. Degano, R. Gorrieri and A. Marchetti-Spaccamela, eds, Proceedings of ICALP ’97, volume 1256 of LNCS, pages 325–335. Springer, 1997.

[Fu99a] Y. Fu. Open Bisimulations on Chi Processes. In J. C. Baeten and S. Mauw, eds, Proceedings of CONCUR ’99, volume 1664 of LNCS, pages 304–319. Springer, Aug. 1999.

[Fu99b] Y. Fu. Variations on Mobile Processes. Theoretical Computer Science, 221:327–368, 1999.

[FY00a] Y. Fu and Z. Yang. Chi Calculus with Mismatch. In C. Palamidessi, ed, Proceedings of CONCUR 2000, volume 1877 of LNCS, pages 596–610. Springer, Aug. 2000.

[FY00b] Y. Fu and Z. Yang. The Ground Congruence for Chi Calculus. In S. Kapoor and S. Prasad, eds, Proceedings of FSTTCS 2000, volume 1974 of LNCS, pages 385–396. Springer, Dec. 2000.

[Gar99] P. Gardner. Closed action calculi. Theoretical Computer Science, 228(1–2):77–103, Oct. 1999.

[Gar00] P. Gardner. From Process Calculi to Process Frameworks. In C. Palamidessi, ed, Proceedings of CONCUR 2000, volume 1877 of LNCS, pages 69–88. Springer, Aug. 2000.

[Gay93] S. J. Gay. A Sort Inference Algorithm for the Polyadic π-Calculus. In Proceedings of POPL ’93. ACM, January 1993.

[Gay00] S. Gay. A Framework for the Formalisation of Pi-Calculus Type Systems in Isabelle/HOL. Technical Report, University of Glasgow, 2000.

[GC99] A. D. Gordon and L. Cardelli. Equational Properties of Mobile Ambients. In W. Thomas, ed, Proceedings of FoSSaCS ’99, volume 1578 of LNCS, pages 212–226. Springer, 1999.

[GH98] A. D. Gordon and P. D. Hankin. A Concurrent Object Calculus: Reduction and Typing. In U. Nestmann and B. C. Pierce, eds, Proceedings of HLCL ’98, volume 16.3 of ENTCS. Elsevier Science Publishers, 1998.

[GHS02] J. C. Godskesen, T. Hildebrandt and V. Sassone. A Calculus of Mobile Resources. In L. Brim, P. Jančar, M. Křetinský and A. Kučera, eds, Proceedings of CONCUR 2002, volume 2421 of LNCS, pages 272–287. Springer, Aug. 2002.

[GJ01] A. D. Gordon and A. Jeffrey. Authenticity by Typing for Security Protocols. volume 2126 of Lecture Notes in Computer Science. Springer-Verlag, 2001.

[Gla92a] J. Glauert. Asynchronous Mobile Processes and Graph Rewriting. In D. Etiemble and J.-C. Syre, eds, Proceedings of PARLE ’92, volume 605 of LNCS, pages 63–80. Springer, 1992.

[Gla92b] J. Glauert. Parallel Implementation of Functional Languages Using Small Processes. In Parallel Implementation Workshop. RWTH Aachen, 1992.

[GLT92] J. Glauert, L. Leth and B. Thomsen. A New Process Model for Functions. In Term Graph Rewriting: Theory and Practice. John Wiley & Sons Ltd, 1992.

[GLW02] P. Gardner, C. Laneve and L. Wischik. The Fusion Machine (Extended Abstract). In L. Brim, P. Jančar, M. Křetinský and A. Kučera, eds, Proceedings of CONCUR 2002, volume 2421 of LNCS, pages 418–433. Springer, Aug. 2002.

[GM96] A. D. Gordon and T. Melham. Five Axioms of Alpha-Conversion. In Proceedings of TPHOL’96, volume 1125 of LNCS, pages 173–190. Springer, 1996.

[GR97a] S. Gnesi and G. Ristori. Model Checking a Mobile Telephone System. In S. Gnesi and D. Latella, eds, Proceedings of FMICS ’97, pages 113–123. Consiglio Nazionale Ricerche di Pisa, 1997.

[GR97b] S. Gnesi and G. Ristori. A Model Checking Algorithm for π-Calculus Agents. In Proceedings of Italian Conference on Theoretical Computer Science ’97. Kluwer Academic Publishers, 1997.

[GW00] P. Gardner and L. Wischik. Explicit fusions. In M. Nielsen and B. Rovan, eds, Proceedings of MFCS 2000, volume 1893 of LNCS, pages 373–382. Springer, 2000.

[Hav94] K. Havelund. The Fork Calculus — A Theory of Communication, Concurrency and Process Creation. PhD thesis, University of Copenhagen, Jan. 1994.

[Hen91] M. Hennessy. A Model for the π Calculus. CogSci Report 8/91, University of Sussex, Brighton, 1991.

[Hen94a] M. Hennessy. A Fully Abstract Denotational Model for Higher-Order Processes. Journal of Information and Computation, 112(1):55–95, 1994. An extended abstract appeared in Proceedings of LICS ’93, pages 397–408.

[Hen94b] M. Hennessy. Higher-Order Processes and Their Models. In S. Abitboul and E. Shamir, eds, Proceedings of ICALP ’94, volume 820 of LNCS, pages 286–303. Springer, 1994.

[Hen96] M. Hennessy. A Fully Abstract Denotational Model for the π-Calculus. CogSci Report 96:04, University of Sussex, Brighton, June 1996.

[Hen99] L. Henry-Gréard. Proof of the Subject Reduction Property for a Pi-Calculus in Coq. Rapport de Recherche RR-3698, INRIA, 1999.

[Hen00] M. Hennessy. The security pi-calculus and non-interference. Computer Science Technical Report 2000:05, School of Cognitive and Computing Sciences, University of Sussex, 2000.

[HHK95] M. Hansen, H. Hüttel and J. Kleist. Bisimulations for asynchronous mobile processes. In I. Lee and S. A. Smolka, eds, Proceedings of CONCUR ’95, volume 962 of LNCS. Springer, 1995.

[Hir94] K. Hirata. Embedding of Moded Flat GHC into Polyadic π-Calculus. In International Symposium on Fifth Generation Computer Systems 1994, Workshop on Parallel Logic Programming, pages 3–16. ICOT, 1994.

[Hir97] D. Hirschkoff. A Full Formalization of Pi-Calculus Theory in the Calculus of Constructions. In E. L. Gunter and A. Felty, eds, Theorem Proving in Higher-Order Logic: 10th International Conference, TPHOLs’97, volume 1275 of LNCS, pages 153–169, 1997.

[Hir98] D. Hirschkoff. Automatically Proving Up to Bisimulation. In P. Jancar and M. Kretinsky, eds, Proceedings of MFCS ’98 Workshop on Concurrency, volume 18 of ENTCS. Elsevier Science Publishers, 1998.

[Hir99] D. Hirschkoff. On the Benefits of Using the Up-To Techniques for Bisimulation Verification. In W. R. Cleaveland, ed, Proceedings of TACAS ’99, volume 1579 of LNCS, pages 286–299. Springer, 1999.

[HJNN99] R. R. Hansen, J. G. Jensen, F. Nielson and H. R. Nielson. Abstract Interpretation of Mobile Ambients. In A. Cortesi and G. Filé, eds, Proceedings of SAS ’99, volume 1694 of LNCS, pages 134–148. Springer, 1999.

[HK96] H. Hüttel and J. Kleist. Objects as Mobile Processes. Research Series RS-96-38, BRICS, Oct. 1996. Presented at MFPS ’96.

[HKNS98] H. Hüttel, J. Kleist, U. Nestmann and D. Sangiorgi. Surrogates in Øjeblik: Towards Migration in Obliq. In H. Hüttel and U. Nestmann, eds, Proceedings of SOAP ’98, volume NS-98-5 of BRICS Notes Series, pages 43–50, 1998.

[HLMP98] F. Honsell, M. Lensia, U. Montanari and M. Pistore. Final Semantics for the π-Calculus. In D. Gries and W.-P. de Roever, eds, Proceedings of PROCOMET ’98, pages 225–243, 1998.

[HMS01] F. Honsell, M. Miculan and I. Scagnetto. π-calculus in (Co)Inductive Type Theory. Theoretical Computer Science, 253(2):239–285, 2001.

[Hon92a] K. Honda. Notes on Soundness of a Mapping from π-calculus to ν-calculus. With comments added in October 1993, May 1992.

[Hon92b] K. Honda. Two Bisimilarities in ν-Calculus. CS report 92-002, Keio University, 1992. Revised on March 31, 1993.

[Hon93] K. Honda. Types for Dyadic Interaction. In E. Best, ed, Proceedings of CONCUR ’93, volume 715 of LNCS, pages 509–523. Springer, 1993.

[Hon96] K. Honda. Composing Processes. In Proceedings of POPL ’96, pages 344–357. ACM, Jan. 1996.

[HP00] O. M. Herescu and C. Palamidessi. Probabilistic Asynchronous π-Calculus. In J. Tiuryn, ed, Proceedings of FoSSaCS 2000, volume 1784 of LNCS, pages 146–160. Springer, 2000.

[HP02] O. M. Herescu and C. Palamidessi. A Randomized Distributed Encoding of the Pi-Calculus with Mixed Choice. In R. A. Baeza-Yates, U. Montanari and N. Santoro, eds, IFIP TCS, volume 223 of IFIP Conference Proceedings, pages 537–549. Kluwer, 2002.

[HR98] M. Hennessy and J. Riely. Resource Access Control in Systems of Mobile Agents. In U. Nestmann and B. C. Pierce, eds, Proceedings of HLCL ’98, volume 16.3 of ENTCS, pages 3–17. Elsevier Science Publishers, 1998. Full version as CogSci Report 2/98, University of Sussex, Brighton.

[HR99] M. Hennessy and J. Riely. Type-safe Execution of Mobile Agents in Anonymous Networks. In J. Vitek and C. D. Jensen, eds, Proceedings of MOS ’98, volume 1603 of LNCS. Springer, 1999. Full version as CogSci Report 3/98, University of Sussex, Brighton.

[HR00a] M. Hennessy and J. Riely. Information flow vs. resource access in the asynchronous pi-calculus. Computer Science Technical Report 2000:03, School of Cognitive and Computing Sciences, University of Sussex, 2000.

[HR00b] M. Hennessy and J. Riely. Information Flow vs. Resource Access in the Information Asynchronous Pi-Calculus. In U. Montanari, J. Rolim and E. Welzl, eds, Proceedings of ICALP 2000, volume 1853 of LNCS, page ? Springer, July 2000.

[HT91a] K. Honda and M. Tokoro. An Object Calculus for Asynchronous Communication. In P. America, ed, Proceedings of ECOOP ’91, volume 512 of LNCS, pages 133–147. Springer, July 1991.

[HT91b] K. Honda and M. Tokoro. A Small Calculus for Concurrent Objects. ACM OOPS Messenger, 2(2):50–54, 1991.

[HT92] K. Honda and M. Tokoro. On Asynchronous Communication Semantics. In M. Tokoro, O. Nierstrasz and P. Wegner, eds, Object-Based Concurrent Computing 1991, volume 612 of LNCS, pages 21–51. Springer, 1992.

[Hüt02] H. Hüttel. Deciding Framed Bisimilarity. In A. Kucera and R. Mayr, eds, Proceedings of INFINITY 2002, volume 68 of ENTCS. Elsevier Science Publishers, 2002.

[HVK98] K. Honda, V. T. Vasconcelos and M. Kubo. Language Primitives and Type Discipline for Structured Communication-Based Programming. In C. Hankin, ed, Proceedings of ESOP ’98, volume 1381 of LNCS, pages 122–138. Springer, 1998.

[HVY00] K. Honda, V. Vasconcelos and N. Yoshida. Secure Information Flow as Typed Process Behaviour. In G. Smolka, ed, Proceedings of ESOP 2000, volume 1782 of LNCS, pages 180–199. Springer, 2000.

[HY94a] K. Honda and N. Yoshida. Combinatory Representation of Mobile Processes. In Proceedings of POPL ’94, pages 348–360. ACM, January 1994.

[HY94b] K. Honda and N. Yoshida. Replication in Concurrent Combinators. In M. Hagiya and J. C. Mitchell, eds, Proceedings of TACS ’94, volume 789 of LNCS, pages 786–805. Springer, 1994.

[HY95] K. Honda and N. Yoshida. On Reduction-Based Process Semantics. Theoretical Computer Science, 152(2):437–486, 1995. An extract appeared in Proceedings of FSTTCS ’93, LNCS 761.

[IK00] A. Igarashi and N. Kobayashi. Type Reconstruction for Linear Pi-Calculus with I/O Subtyping. Journal of Information and Computation, 161(1):1–44, 2000. An extended abstract appeared in the Proceedings of SAS ’97, LNCS 1302.

[IK01] A. Igarashi and N. Kobayashi. A Generic Type System for the Pi-Calculus. In Proceedings of POPL ’01, pages 128–141. ACM, Jan. 2001.

[JJ95] L. J. Jagadeesan and R. Jagadeesan. Causality and True Concurrency: A Data-flow Analysis of the Pi-Calculus. In V. S. Alagar and M. Nivat, eds, Proceedings of AMAST ’95, volume 936 of LNCS, pages 277–291. Springer, 1995.

[Jon93a] C. Jones. A Pi-Calculus Semantics for an Object-Based Design Notation. In E. Best, ed, Proceedings of CONCUR ’93, volume 715 of LNCS, pages 158–172. Springer, 1993.

[Jon93b] C. Jones. Process Algebra Foundations for an Object-Based Design Notation. Technical Report UMCS-93-10-1, University of Manchester, Oct. 1993.

[Jon94] C. Jones. Process Algebra Arguments about an Object-Based Design Notation. In A. W. Roscoe, ed, A Classical Mind: Essays in Honour of C.A.R. Hoare, pages 1–15. Prentice Hall, 1994.

[Jon95] C. B. Jones. Fixing the semantics of some concurrent object-oriented concepts. In S. Brookes, M. Main, A. Melton and M. Mislove, eds, Proceedings of MFPS ’95, volume 1 of ENTCS. Elsevier Science Publishers, 1995.

[JP90] R. Jagadeesan and P. Panangaden. A Domain-Theoretic Model for a Higher-order Process Calculus. In M. S. Paterson, ed, Proceedings of ICALP ’90, volume 443 of LNCS, pages 181–194. Springer, 1990. Extended version as Technical Report 89-1058, Cornell University.

[JR99] A. Jeffrey and J. Rathke. Towards a theory of bisimulation for local names. In Proceedings of LICS ’99, pages 56–66. IEEE, Computer Society Press, July 1999.

[KH94] J. Kleist and M. Hansen. Process Calculi with Asynchronous Communication. Master’s Thesis, Aalborg University, Aug. 1994. Available as Report R-94-2028, ISSN 0908-1216.

[Kle00] J. Kleist. Reasoning about Objects using Process Calculus Techniques. PhD thesis, Aalborg University, Denmark, May 2000.

[Kob98] N. Kobayashi. A Partially Deadlock-Free Typed Process Calculus. ACM Transactions on Programming Languages and Systems, 20(2):436–482, 1998. An extended abstract previously appeared in the Proceedings of LICS ’97, pages 128–139.

[Kob00] N. Kobayashi. Type Systems for Concurrent Processes: From Deadlock-Freedom to Livelock-Freedom, Time-Boundedness. In J. van Leeuwen, O. Watanabe, M. Hagiya, P. Mosses and T. Ito, eds, Proceedings of TCS 2000, volume 1872 of LNCS, pages 365–389. IFIP, Springer, Aug. 2000.

[Kön99] B. König. Description and Verification of Mobile Processes with Graph Rewriting Techniques. PhD thesis, Technische Universität München, 1999.

[Kön00a] B. König. Analysing Input/Output-Capabilities of Mobile Processes with a Generic Type System. In U. Montanari, J. Rolim and E. Welzl, eds, Proceedings of ICALP 2000, volume 1853 of LNCS, pages 403–414. Springer, July 2000.

[Kön00b] B. König. A General Framework for Types in Graph Rewriting. In S. Kapoor and S. Prasad, eds, Proceedings of FSTTCS 2000, volume 1974 of LNCS, pages 373–384. Springer, Dec. 2000.

[Kön00c] B. König. A Graph Rewriting Semantics for the Polyadic Pi-Calculus. In Workshop on Graph Transformation and Visual Modeling Techniques (Geneva, Switzerland), ICALP Workshops 2000, pages 451–458. Carleton Scientific, 2000.

[KPT96] N. Kobayashi, B. C. Pierce and D. N. Turner. Linearity and the Pi-Calculus. In Proceedings of POPL ’96, pages 358–371. ACM, Jan. 1996.

[KS82] J. Kennaway and M. Sleep. Expressions as Processes. In Symposium on lisp and Functional Programming, pages 21–28. ACM, Aug. 1982.

[KS85] J. Kennaway and M. Sleep. Syntax and Informal Semantics of DyNe, a Parallel Language. In B. T. Denvir, W. T. Harwood, M. I. Jackson and M. J. Wray, eds, The Analysis of Concurrent Systems 1983, volume 207 of LNCS, pages 222–230. Springer, 1985.

[KS98] J. Kleist and D. Sangiorgi. Imperative Objects and Mobile Processes. In D. Gries and W.-P. de Roever, eds, Proceedings of PROCOMET ’98, pages 285–303, 1998.

[KSS00] N. Kobayashi, S. Saito and E. Sumii. An Implicitly-Typed Deadlock-Free Process Calculus. In C. Palamidessi, ed, Proceedings of CONCUR 2000, volume 1877 of LNCS, pages 489–503. Springer, Aug. 2000.

[Lan96] C. Laneve. May and Must Testing in the Join-Calculus. Technical Report UBLCS-96-4, DoCS, University of Bologna, March 1996. Revised: May 1996.

[Lav93] C. Lavatelli. Non-deterministic Lazy λ-calculus vs. π-calculus. Technical Report LIENS-93-15, Ecole Normale Supérieure, Paris, Sept. 1993.

[LC98] Z. Li and H. Chen. Checking Strong/Weak Bisimulation Equivalences and Observation Congruence for the Pi-calculus. In K. G. Larsen, S. Skyum and G. Winskel, eds, Proceedings of ICALP ’98, volume 1443 of LNCS, pages 707–718. Springer, July 1998.

[Let91] L. Leth. Functional Programs as Reconfigurable Networks of Communicating Processes. PhD thesis, Imperial College, University of London, Feb. 1991.

[Let92] L. Leth. A New Direction in Functions as Processes. Technical Report ECRC-92-5, European Computer-Industry Research Centre, Munich, 1992.

[Lév97] J.-J. Lévy. Some Results on the Join-Calculus. In M. Abadi and T. Ito, eds, Proceedings of TACS ’97, volume 1281 of LNCS. Springer, 1997.

[Li94] B. Z. Li. A π-Calculus Specification of Prolog. In D. Sannella, ed, Proceedings of ESOP ’94, volume 788 of LNCS, pages 379–393. Springer, 1994.

[Lin94] H. Lin. Symbolic Bisimulation and Proof Systems for the π-Calculus. Technical Report 7/94, School of Cognitive and Computing Sciences, University of Sussex, UK, 1994.

[Lin95a] H. Lin. Complete Inference Systems for Weak Bisimulation Equivalences in the π-Calculus. In P. D. Mosses, M. Nielsen and M. I. Schwarzbach, eds, Proceedings of TAPSOFT ’95, volume 915 of LNCS, pages 187–201. Springer, 1995. Presented in the CAAP-section. Available as Technical Report ISCAS-LCS-94-11, Institute of Software, Chinese Academy of Sciences, 1994.

[Lin95b] H. Lin. Unique Fixpoint Induction for Mobile Processes. In I. Lee and S. A. Smolka, eds, Proceedings of CONCUR ’95, volume 962 of LNCS, pages 88–102. Springer, 1995.

[Lin98] H. Lin. Complete Proof Systems for Observation Congruences in Finite-Control Picalculus. In K. G. Larsen, S. Skyum and G. Winskel, eds, Proceedings of ICALP ’98, volume 1443 of LNCS, pages 443–454. Springer, July 1998.

[Liu94] X. Liu. Characterizing Bisimulation Congruence in the π-Calculus. In B. Jonsson and J. Parrow, eds, Proceedings of CONCUR ’94, volume 836 of LNCS, pages 331–350. Springer, Aug. 1994. Extended version as CogSci Report 2/94, University of Sussex, Brighton.

[LM92] C. Laneve and U. Montanari. Mobility in the CC-Paradigm. In I. M. Havel and V. Koubek, eds, Proceedings of MFCS ’92, volume 629 of LNCS, pages 336–345. Springer, 1992.

[LM98] F. Le Fessant and L. Maranget. Compiling Join-Patterns. In U. Nestmann and B. C. Pierce, eds, Proceedings of HLCL ’98, volume 16.3 of ENTCS. Elsevier Science Publishers, 1998.

[LM01] F. Levi and S. Maffeis. An Abstract Interpretation Framework for Analysing Mobile Ambients. In Proceedings of SAS ’01, LNCS. Springer, 2001. To appear.

[LS00] F. Levi and D. Sangiorgi. Controlling Interference in Ambients. In Proceedings of POPL ’00, pages 352–364. ACM, Jan. 2000.

[LSN97] M. Lumpe, J.-G. Schneider and O. Nierstrasz. Using Metaobjects to Model Concurrent Objects with PICT. In Proceedings of Langages et Modèles à Objects (Leysin, October 1996), pages 1–12, Oct. 1997.

[LV99] C. Laneve and B. Victor. Solos in Concert. In J. Wiederman, P. van Emde Boas and M. Nielsen, eds, Proceedings of ICALP ’99, volume 1644 of LNCS, pages 513–523. Springer, July 1999.

[LW95a] X. Liu and D. Walker. Confluence of Processes and Systems of Objects. In P. D. Mosses, M. Nielsen and M. I. Schwarzbach, eds, Proceedings of TAPSOFT ’95, volume 915 of LNCS, pages 217–231. Springer, 1995. Presented in the CAAP-section. Available as University of Warwick Research Report CS-RR-272, October 1994.

[LW95b] X. Liu and D. Walker. A Polymorphic Type System for the Polyadic π-Calculus. In I. Lee and S. A. Smolka, eds, Proceedings of CONCUR ’95, volume 962 of LNCS, pages 103–116. Springer, 1995.

[LW98] X. Liu and D. Walker. Partial Confluence of Processes and Systems of Objects. Theoretical Computer Science, 1998.

[LW00] X. Liu and D. Walker. Concurrent Objects as Mobile Processes. In G. Plotkin, C. Stirling and M. Tofte, eds, Proof, Language and Interaction: Essays in Honour of Robin Milner, Foundations of Computing, May 2000.

[Mar02] F. Martinelli. About Compositional Analysis of Pi-Calculus Processes. In R. A. Baeza-Yates, U. Montanari and N. Santoro, eds, IFIP TCS, volume 223 of IFIP Conference Proceedings, pages 524–536. Kluwer, 2002.

[Mel94] T. F. Melham. A Mechanized Theory of the π-Calculus in HOL. Nordic Journal of Computing, 1(1):50–76, 1994.

[Mer98] M. Merro. On the Expressiveness of Chi, Update, and Fusion calculi. In C. Palamidessi and I. Castellani, eds, Proceedings of EXPRESS ’98, volume 16.2 of ENTCS. Elsevier Science Publishers, 1998.

[Mer99] M. Merro. On Equators in Asynchronous Name-passing Calculi without Matching. In I. Castellani and B. Victor, eds, Proceedings of EXPRESS ’99, volume 27 of ENTCS. Elsevier Science Publishers, 1999.

[Mer00a] M. Merro. Locality and Polyadicity in Asynchronous Name-Passing Calculi. In J. Tiuryn, ed, Proceedings of FoSSaCS 2000, volume 1784 of LNCS, pages 238–251. Springer, 2000.

[Mer00b] M. Merro. Locality in the π-calculus and applications to distributed objects. PhD thesis, Ecole des Mines, France, October 2000.

[Mil91] R. Milner. Sorts in the π-Calculus (Extended Abstract). In E. Best and G. Rozenberg, eds, Proceedings of the 3rd Workshop on Concurrency and Compositionality, volume 191 of GMD-Studien. GMD Bonn, St. Augustin, 1991. Also available as Report 6/91 from University of Hildesheim.

[Mil92a] R. Milner. Action Structures. Technical Report ECS-LFCS-92-249, LFCS, University of Edinburgh, 1992.

[Mil92b] R. Milner. Action Structures and Polynomial π-Calculus. In E. Best, ed, What Good are Partial Orders?, Proceedings of a CALIBAN-Workshop, Sheffield, pages 77–105. Universität Hildesheim, Oct. 1992. Hildesheimer Informatik-Berichte 13/92.

[Mil92c] R. Milner. Functions as Processes. Journal of Mathematical Structures in Computer Science, 2(2):119–141, 1992. Previous version as Rapport de Recherche 1154, INRIA Sophia-Antipolis, 1990, and in Proceedings of ICALP ’91, LNCS 443.

[Mil93a] D. Miller. The π-calculus as a Theory in Linear Logic. In E. Lamma and P. Mello, eds, Proceedings of the 1992 Workshop on Extensions to Logic Programming, volume 660 of Lecture Notes in Computer Science, pages 242–265, 1993.

[Mil93b] R. Milner. Action Calculi, or Concrete Action Structures. In A. M. Borzyszkowski and S. Sokołowski, eds, Proceedings of MFCS ’93, volume 711 of LNCS, pages 105–121. Springer, Sept. 1993. Also titled Action Calculi I: Axioms and Applications.

[Mil93c] R. Milner. An Action Structure for the Synchronous π-Calculus. In Z. Ésik, ed, Proceedings of FCT’93, volume 710 of LNCS, pages 87–105. Springer, 1993.

[Mil93d] R. Milner. Elements of Interaction. Communications of the ACM, 36(1):78–89, 1993. Turing Award Lecture.

[Mil93e] R. Milner. The Polyadic π-Calculus: A Tutorial. In F. L. Bauer, W. Brauer and H. Schwichtenberg, eds, Logic and Algebra of Specification, volume 94 of Series F. NATO ASI, Springer, 1993. Available as Technical Report ECS-LFCS-91-180, University of Edinburgh, October 1991.

[Mil93f] R. Milner. Action Structures for the π-Calculus. Technical Report ECS-LFCS-93-264, LFCS, University of Edinburgh, May 1993.

[Mil94] R. Milner. Pi-Nets: A Graphical Form of π-Calculus. In D. Sannella, ed, Proceedings of ESOP ’94, volume 788 of LNCS, pages 26–42. Springer, 1994.

[Mil95] R. Milner. Action Structures and the π-Calculus. In H. Schwichtenberg, ed, Proof and Computation, volume 139 of Series F. NATO ASI, Springer, 1995. Available from

ftp://ftp.cl.cam.ac.uk/users/rm135/ap.ps.Z.

[Mil96] R. Milner. Calculi for Interaction. Acta Informatica, 33(8):707–737, Nov. 1996.

[Mil99] R. Milner. Communicating and Mobile Systems: the π-Calculus. Cambridge University Press, May 1999.

[Mit01] J. C. Mitchell. Probabilistic Polynomial-Time Process Calculus and Security Protocol Analysis. In D. Sands, ed, Proceedings of ESOP 2001, volume 2028 of LNCS, pages 23–29. Springer, 2001.

[MKN00] M. Merro, J. Kleist and U. Nestmann. Local π-Calculus at Work: Mobile Objects as Mobile Processes. In J. van Leeuwen, O. Watanabe, M. Hagiya, P. Mosses and T. Ito, eds, Proceedings of TCS 2000, volume 1872 of LNCS, pages 390–408. IFIP, Springer,

Aug. 2000. Available from http://www.cs.auc.dk/research/FS/ojeblik/. Full ver-

sion accepted for publication in Journal of Information and Computation.

[MP95a] U. Montanari and M. Pistore. Checking Bisimilarity for Finitary π-Calculus. In I. Lee and S. A. Smolka, eds, Proceedings of CONCUR ’95, volume 962 of LNCS, pages 42–56. Springer, 1995.

[MP95b] U. Montanari and M. Pistore. Concurrent Semantics for the π-calculus. In S. Brookes, M. Main, A. Melton and M. Mislove, eds, Proceedings of MFPS ’95, volume 1 of ENTCS. Elsevier Science Publishers, 1995.

[MP97] U. Montanari and M. Pistore. An Introduction to History Dependent Automata. In A. Gordon, A. Pitts and C. Talcott, eds, Conference Record of HOOTS II, volume 10 of ENTCS. Elsevier Science Publishers, 1997.

[MP99] U. Montanari and M. Pistore. Finite State Verification for the Asynchronous Pi-Calculus. In W. R. Cleaveland, ed, Proceedings of TACAS ’99, volume 1579 of LNCS, pages 255–269. Springer, 1999.

[MPW92] R. Milner, J. Parrow and D. Walker. A Calculus of Mobile Processes, Part I/II. Journal of Information and Computation, 100:1–77, Sept. 1992.

[MPW93] R. Milner, J. Parrow and D. Walker. Modal logics for mobile processes. Theoretical Computer Science, 114:149–171, 1993.

[MS92] R. Milner and D. Sangiorgi. Barbed Bisimulation. In W. Kuich, ed, Proceedings of ICALP ’92, volume 623 of LNCS, pages 685–695. Springer, 1992.

[MS98] M. Merro and D. Sangiorgi. On Asynchrony in Name-Passing Calculi. In K. G. Larsen, S. Skyum and G. Winskel, eds, Proceedings of ICALP ’98, volume 1443 of LNCS, pages 856–867. Springer, July 1998.

[MS02] M. Merro and V. Sassone. Typing and Subtyping Mobility in Boxed Ambients. In L. Brim, P. Jančar, M. Křetinský and A. Kučera, eds, Proceedings of CONCUR 2002, volume 2421 of LNCS, pages 306–329. Springer, Aug. 2002.

[MT97] U. Montanari and C. Talcott. Can Actors and π-Agents Live Together. In A. Gordon, A. Pitts and C. Talcott, eds, Conference Record of HOOTS II, volume 10 of ENTCS. Elsevier Science Publishers, 1997.

[Nes96] U. Nestmann. On Determinacy and Nondeterminacy in Concurrent Programming. PhD thesis, Technische Fakultät, Universität Erlangen, November 1996. Arbeitsbericht IMMD-29(14).

[Nes98] U. Nestmann. On the Expressive Power of Joint Input. In C. Palamidessi and I. Castellani, eds, Proceedings of EXPRESS ’98, volume 16.2 of ENTCS. Elsevier Science Publishers, 1998.

[Nes00] U. Nestmann. What Is a ‘Good’ Encoding of Guarded Choice? Journal of Information and Computation, 156:287–319, 2000. An extended abstract appeared in the Proceedings of EXPRESS ’97, volume 7 of ENTCS.

[Nie96] J. Niehren. Functional Computation as Concurrent Computation. In Proceedings of POPL ’96, pages 333–343. ACM, Jan. 1996. Full version as Research Report RR-95-14, DFKI Saarbrücken.

[NN00] H. R. Nielson and F. Nielson. Shape Analysis for Mobile Ambients. In Proceedings of POPL ’00, pages 142–154. ACM, Jan. 2000.

[NNHJ99] F. Nielson, H. R. Nielson, R. R. Hansen and J. G. Jensen. Validating Firewalls in Mobile Ambients. In J. C. Baeten and S. Mauw, eds, Proceedings of CONCUR ’99, volume 1664 of LNCS, pages 463–477. Springer, Aug. 1999.

[NNS00] F. Nielson, H. R. Nielson and M. Sagiv. A Kleene Analysis of Mobile Ambients. In G. Smolka, ed, Proceedings of ESOP 2000, volume 1782 of LNCS, pages 305–319. Springer, 2000.

[NP00] U. Nestmann and B. C. Pierce. Decoding Choice Encodings. Journal of Information and Computation, 163:1–59, 2000. Also available as report BRICS-RS-99-42, Universities of Aalborg and Århus, Denmark, 1999. An extended abstract appeared in the Proceedings of CONCGUR ’96, LNCS 1119, pages 179–194.

[NR99] U. Nestmann and A. Ravara. Semantics of Objects As Processes (SOAP). In A. Moreira and S. Demeyer, eds, ECOOP ’99 Workshop Reader, volume 1743 of LNCS, pages 314–325. Springer, Oct. 1999. An introduction to, and summary of, the 2nd International SOAP-Workshop.

[NS96] U. Nestmann and M. Steffen. Correct Transformational Design of Concurrent Search Structures. In B. Freitag, C. Jones, C. Lengauer and H.-J. Schek, eds, Object-Orientation with Parallelism and Persistence, pages 23–42. Kluwer Academic Publishers, 1996.

[NS97] U. Nestmann and M. Steffen. Typing Confluence. In S. Gnesi and D. Latella, eds, Proceedings of FMICS ’97, pages 77–101. Consiglio Nazionale Ricerche di Pisa, 1997. Also available as report ERCIM-10/97-R052, European Research Consortium for Informatics and Mathematics, 1997.

[NV98] U. Nestmann and B. Victor. Calculi for Mobile Processes: Bibliography and Web Pages. Bulletin of the EATCS, 64:139–144, Feb. 1998. Contribution to the Concurrency Column.

[NW02] M. Nygaard and G. Winskel. HOPLA: A Higher-Order Process Language. In L. Brim, P. Jančar, M. Křetinský and A. Kučera, eds, Proceedings of CONCUR 2002, volume 2421 of LNCS, pages 434–448. Springer, Aug. 2002.

[Ode95a] M. Odersky. Applying π: Towards a Basis for Concurrent Imperative Programming. In U. S. Reddy, ed, Proceedings of SIPL ’95, pages 95–108. ACM/SIGPLAN, University of Illinois at Urbana-Champaign, January 22 1995. Available as Technical Report UIUCDCS-R-95-1900.

[Ode95b] M. Odersky. Polarized Name Passing. In P. S. Thiagarajan, ed, Proceedings of FSTTCS ’95, volume 1026 of LNCS, pages 324–337. Springer, 1995.

[Ode00] M. Odersky. Functional Nets. In G. Smolka, ed, Proceedings of ESOP 2000, volume 1782 of LNCS, pages 1–25. Springer, 2000.

[Oka93] M. Okada. Mobile Linear Logic as a Framework for Asynchronous and Synchronous Mobile Communication Calculi (Preliminary Report). Concordia University, Mar. 1993.

[OP92] F. Orava and J. Parrow. An Algebraic Verification of a Mobile Network. Journal of Formal Aspects of Computing, 4:497–543, 1992. Missing figures available at http:

//www.it.kth.se/~joachim/handfig.ps.gz.

[Ora94] F. Orava. On the Formal Analysis of Telecommunication Protocols. PhD thesis, Department of Computer Systems, Uppsala University, Sweden, May 1994.

[Pal97] C. Palamidessi. Comparing the Expressive Power of the Synchronous and the Asynchronous π-calculus. In Proceedings of POPL ’97, pages 256–265. ACM, Jan. 1997.

[Par95] J. Parrow. Interaction Diagrams. Nordic Journal of Computing, 2:407–443, 1995. A previous version appeared in Proceedings of A Decade in Concurrency, LNCS 803: 477–508, 1993.

[Par99] J. Parrow. On the relationship between two proof systems for the pi-calculus. Un-

published, available from http://www.it.kth.se/~joachim/compl.txt, 1999.

[Par00] J. Parrow. Trios in Concert. In G. Plotkin, C. Stirling and M. Tofte, eds, Proof, Language and Interaction: Essays in Honour of Robin Milner, Foundations of Computing, pages 621–637, May 2000.

[Par01] J. Parrow. An Introduction to the pi-Calculus. In J. Bergstra, A. Ponse and S. Smolka, eds, Handbook of Process Algebra, pages 479–543. Elsevier Science, 2001.

[Pav97] D. Pavlovic. Categorical Logic of Names and Abstraction in Action Calculi. Mathematical Structures in Computer Science, 7(6):619–637, 1997.

[Phi96] A. Philippou. Reasoning about Systems with Evolving Structure. PhD thesis, University of Warwick, Dec. 1996.

[Pie97] B. C. Pierce. Foundational Calculi for Programming Languages. In A. B. Tucker, ed, Handbook of Computer Science and Engineering, chapter 139, pages 2190–2207. CRC Press, 1997.

[Pri95] C. Priami. Stochastic π-Calculus. The Computer Journal, 38(6):578–589, 1995. Proceedings of PAPM ’95, the 3rd Workshop on Process Algebras and Performance Modelling.

[Pri96a] C. Priami. Enhanced Operational Semantics for Concurrency. PhD thesis, Dipartimento di Informatica, Università di Pisa-Genova-Udine, Mar. 1996. Available as PhD number TD-08/96.

[Pri96b] C. Priami. Integrating Behavioural and Performance Analysis with Topology Information. In Proceedings of 29^{th} Hawaiian International Conference on System Sciences, volume 1, pages 508–516, Maui, Hawaii, 1996. IEEE.

[Pri96c] C. Priami. Interleaving-based Partial Ordering Semantics. In Proceedings of Italian Conference on Theoretical Computer Science, pages 264–278, Ravello, November 1995, 1996. World Scientific.

[Pri96d] C. Priami. Stochastic pi-calculus with General Distributions. In CLUP, ed, Proceedings of PAPM ’96, 1996.

[Pri97] C. Priami. Stochastic Analysis of Mobile Telephony Networks. In E. Brinskma and A. Nymeyer, eds, Proceedings of PAPM ’97, pages 145–171, 1997.

[PRT93] B. C. Pierce, D. Rémy and D. N. Turner. A Typed Higher-Order Programming Language Based on the Pi-Calculus. In Workshop on Type Theory and its Application to Computer Systems, Kyoto University, July 1993.

[PS95] J. Parrow and D. Sangiorgi. Algebraic Theories for Name-Passing Calculi. Journal of Information and Computation, 120(2):174–197, 1995. A previous version appeared in Proceedings of A Decade in Concurrency, LNCS 803: 477–508, 1993.

[PS96a] B. C. Pierce and D. Sangiorgi. Typing and Subtyping for Mobile Processes. Mathematical Structures in Computer Science, 6(5):409–454, 1996. An extract appeared in Proceedings of LICS ’93: 376–385.

[PS96b] M. Pistore and D. Sangiorgi. A Partition Refinement Algorithm for the π-Calculus. In R. Alur, ed, Proceedings of CAV ’96, volume 1102 of LNCS, pages 38–49, 1996. Full version to appear in Journal of Information and Computation.

[PS00] B. C. Pierce and D. Sangiorgi. Behavioral Equivalence in the Polymorphic Pi-Calculus. Journal of the ACM, 47(3):531–584, 2000. An extended abstract appeared in Proceedings of POPL ’97 : 241–255. Full version available as University of Pennsylvania Technical Report MS-CIS-99-10.

[PT95] B. C. Pierce and D. N. Turner. Concurrent Objects in a Process Calculus. In T. Ito and A. Yonezawa, eds, Proceedings of TPPP ’94, volume 907 of LNCS, pages 187–215. Springer, 1995.

[PT00] B. C. Pierce and D. N. Turner. Pict: A Programming Language Based on the Pi-Calculus. In G. Plotkin, C. Stirling and M. Tofte, eds, Proof, Language and Interaction: Essays in Honour of Robin Milner, Foundations of Computing, May 2000.

[PV97] J. Parrow and B. Victor. The Update Calculus. In M. Johnson, ed, Proceedings of AMAST ’97, volume 1349 of LNCS, pages 409–423. Springer, Dec. 1997. Full version available as Technical report DoCS 97/93, Uppsala University.

[PV98a] J. Parrow and B. Victor. The Fusion Calculus: Expressiveness and Symmetry in Mobile Processes. In Proceedings of LICS ’98, pages 176–185. IEEE, Computer Society Press, July 1998.

[PV98b] J. Parrow and B. Victor. The Tau-Laws of Fusion. In D. Sangiorgi and R. de Simone, eds, Proceedings of CONCUR ’98, volume 1466 of LNCS, pages 99–114. Springer, Sept. 1998.

[PW95] A. Philippou and D. Walker. On Sharing and Determinacy in Concurrent Systems. In I. Lee and S. A. Smolka, eds, Proceedings of CONCUR ’95, volume 962 of LNCS, pages 456–470. Springer, 1995.

[PW97a] A. Philippou and D. Walker. On confluence in the π-calculus. In P. Degano, R. Gorrieri and A. Marchetti-Spaccamela, eds, Proceedings of ICALP ’97, volume 1256 of LNCS, pages 314–324. Springer, 1997.

[PW97b] A. Philippou and D. Walker. A Rigorous Analysis of Concurrent Operations on B-Trees. In A. Mazurkiewicz and J. Winkowski, eds, Proceedings of CONCUR ’97, volume 1243 of LNCS, pages 361–375. Springer, 1997.

[PW97c] A. Philippou and D. Walker. Social Confluence in Client-Server Systems. In Proceedings of CSL ’96, volume 1258 of LNCS, pages 385–398. Springer, 1997.

[PW98] A. Philippou and D. Walker. On Transformations of Concurrent Object Programs. Theoretical Computer Science, 195(2):259–289, 1998. An extended abstract appeared in Proceedings of CONCUR ’96, LNCS 1119: 131–146.

[Qua96] P. Quaglia. The π-Calculus with Explicit Substitutions. PhD thesis, Dipartimento di Informatica, Università di Pisa-Genova-Udine, Mar. 1996. Available as PhD number TD-09/96.

[Qua97] P. Quaglia. On the Finitary Characterization of pi-Congruences. Report Series RS-97-52, BRICS, Denmark, Dec. 1997.

[Qua99] P. Quaglia. Pi-Congruences as CCS Equivalences. In A. M. Haeberer, ed, Proceedings of AMAST ’98, volume 1548 of LNCS, pages 358–372. Springer, 1999.

[QW98] P. Quaglia and D. Walker. On encoding pπ in mπ. In V. Arvind and R. Ramanujam, eds, Proceedings of FSTTCS ’98, volume 1530 of LNCS, pages 42–51. Springer, Dec. 1998.

[QW00] P. Quaglia and D. Walker. On Synchronous and Asynchronous Mobile Processes. In J. Tiuryn, ed, Proceedings of FoSSaCS 2000, volume 1784 of LNCS, pages 283–296. Springer, 2000.

[RE94a] M. Radestock and S. Eisenbach. Towards a Minimal Object-Oriented Language for Distributed and Concurrent Programming. In Thirteenth Annual Symposium on Principles of Distributed Computing (PODC), page 383. ACM, 1994. Abstract at the Poster Session. Full version as Technical Report, Imperial College, University of London.

[RE94b] M. Radestock and S. Eisenbach. What Do You Get From a π-calculus Semantics? In C. Halatsis, D. Maritsas, G. Philokyprou and S. Theodoridis, eds, Proceedings of PARLE ’94, volume 817 of LNCS, pages 635–647. Springer, 1994.

[RH97] J. Riely and M. Hennessy. Distributed Processes and Location Failures. In P. Degano, R. Gorrieri and A. Marchetti-Spaccamela, eds, Proceedings of ICALP ’97, volume 1256 of LNCS, pages 471–481. Springer, 1997. Full version as CogSci Report 2/97, University of Sussex, Brighton.

[RH98] J. Riely and M. Hennessy. A Typed Language for Distributed Mobile Processes. In Proceedings of POPL ’98. ACM, 1998.

[RH99] J. Riely and M. Hennessy. Trust and Partial Typing in Open Systems of Mobile Agents. In Proceedings of POPL ’99, pages 93–104. ACM, 1999. Full version as CogSci Report 4/98, University of Sussex, Brighton.

[RH01] J. Riely and M. Hennessy. Distributed Processes and Location Failures. Theoretical Computer Science, 226:693–735, 2001.

[RHB01] C. Röckl, D. Hirschkoff and S. Berghofer. Higher-Order Abstract Syntax with Induction in Isabelle/HOL: Formalizing the Pi-Calculus and Mechanizing the Theory of Contexts. In F. Honsell and M. Miculan, eds, Proceedings of FoSSaCS 2001, volume 2030 of LNCS, pages 364–378. Springer, 2001.

[RL99] A. Ravara and L. Lopes. Programming and Implementation Issues in Non-Uniform TyCO. Technical Report DCC-99-1, Universidade do Porto, 1999.

[Ros93] B. Ross. A π-calculus semantics of logical variables and unification. In S. Purushothaman and A. Zwarico, eds, Proceedings of First North American Process Algebra Workshop, Stony Brook, 1992, Workshops in Computing, pages 216–230. Springer, 1993.

[RPJ97] B. Robben, F. Piessens and W. Joosen. Formalizing Correlate — from Practice to Pi. In Proceedings of the 2nd BCS-FACS Northern Formal Methods Workshop. Springer, 1997.

[RR01] S. K. Rajamani and J. Rehof. A Behavioral Module System for the Pi-Calculus. In Proceedings of SAS ’01, LNCS. Springer, 2001. To appear.

[RRV98] A. Ravara, P. Resende and V. Vasconcelos. Towards an Algebra of Dynamic Object Types. In H. Hüttel and U. Nestmann, eds, Proceedings of SOAP ’98, volume NS-98-5 of BRICS Notes Series, pages 25–30, 1998.

[RS95] N. Raja and R. K. Shyamasundar. Type Systems for Mobile Processes. In S. Bai, J. Fan and X. Li, eds, The Fourth International Conference for Young Computer Scientists (ICYCS’95), pages 622–629, Beijing, P. R. China, July 1995. Peking University Press.

[RS96] N. Raja and R. K. Shyamasundar. Actors as a Coordinating Model of Computation. In D. Bjorner, M. Broy and I. V. Pottosin, eds, Proceedings of PSI ’96, volume 1181 of LNCS, pages 191–202. Springer, 1996.

[RS97] N. Raja and R. K. Shyamasundar. Combinatory Formulations of Concurrent Languages. ACM Transactions on Programming Languages and Systems, 19(6):899–915, 1997. An extract appeared in Proceedings of ACSC ’95, LNCS 1023: 156–170.

[RS99] C. Röckl and D. Sangiorgi. A π-calculus Process Semantics of Concurrent Idealised ALGOL. In W. Thomas, ed, Proceedings of FoSSaCS ’99, volume 1578 of LNCS, pages 306–322. Springer, 1999.

[RV97] A. Ravara and V. T. Vasconcelos. Behavioural Types for a Calculus of Concurrent Objects. In C. Lengauer, M. Griebl and S. Gorlatch, eds, Proceedings of 3rd International EURO-PAR Conference, volume 1300 of LNCS, pages 554–561. Springer, 1997. Preprint as Technical Report 6/97 from Departamento de Matemática, Instituto Superior Técnico, Universidade Técnica de Lisboa.

[RV00] A. Ravara and V. T. Vasconcelos. Typing Non-Uniform Concurrent Objects. In C. Palamidessi, ed, Proceedings of CONCUR 2000, volume 1877 of LNCS, pages 474– 488. Springer, Aug. 2000.

[San93a] D. Sangiorgi. Expressing Mobility in Process Algebras: First-Order and Higher-Order Paradigms. PhD thesis, LFCS, University of Edinburgh, 1993. CST-99-93 (also published as ECS-LFCS-93-266).

[San93b] D. Sangiorgi. From π-calculus to Higher-Order π-calculus — and back. In M.-C. Gaudel and J.-P. Jouannaud, eds, Proceedings of TAPSOFT ’93, volume 668 of LNCS, pages 151–166. Springer, 1993.

[San94a] D. Sangiorgi. An investigation into functions as processes. In S. Brookes, M. Main, A. Melton and D. Schmidt, eds, Proceedings of MFPS ’93, volume 802 of LNCS, pages 143–159. Springer, 1994.

[San94b] D. Sangiorgi. The Lazy Lambda Calculus in a Concurrency Scenario. Journal of Information and Computation, 111(1):120–131, 1994. An extended abstract appeared in Proceedings of LICS ’92.

[San95] D. Sangiorgi. Lazy functions and mobile processes. Rapport de Recherche RR-2515, INRIA Sophia-Antipolis, 1995.

[San96a] D. Sangiorgi. Bisimulation in Higher-Order Process Calculi. Journal of Information and Computation, 131:141–178, 1996. Available as Rapport de Recherche RR-2508, INRIA Sophia-Antipolis, 1995. An early version appeared in Proceedings of PRO-COMET’94, pages 207–224. IFIP. North Holland Publisher.

[San96b] D. Sangiorgi. Locality and non-interleaving semantics in calculi for mobile processes. Theoretical Computer Science, 155, 1996. See also Report ECS-LFCS-94-282, University of Edinburgh. An extract appeared in Proceedings of TACS’94, LNCS 789.

[San96c] D. Sangiorgi. π-Calculus, Internal Mobility and Agent-Passing Calculi. Theoretical Computer Science, 167(1,2):235–274, 1996. Also as Rapport de Recherche RR-2539, INRIA Sophia-Antipolis, 1995. Extracts of parts of the material contained in this paper can be found in Proceedings of TAPSOFT ’95 and ICALP ’95.

[San96d] D. Sangiorgi. A Theory of Bisimulation for the π-calculus. Acta Informatica, 33:69–97, 1996. Earlier version published as Report ECS-LFCS-93-270, University of Edinburgh. An extended abstract appeared in the Proceedings of CONCUR ’93, LNCS 715.

[San98a] D. Sangiorgi. An Interpretation of Typed Objects into Typed π-Calculus. 143(1):34– 73, 1998. Earlier version published as Rapport de Recherche RR-3000, INRIA Sophia-Antipolis, August 1996.

[San98b] D. Sangiorgi. Interpreting Functions as pi-Calculus Processes: a Tutorial. Rapport de Recherche RR-3470, INRIA Sophia-Antipolis, Aug. 1998.

[San98c] D. Sangiorgi. On the bisimulation proof method. Mathematical Structures in Computer Science, 8(5):447–479, 1998. An extended abstract appeared in the Proceedings of MFCS ’95, LNCS 969: 479–488.

[San99a] D. Sangiorgi. From λ to π; or, Rediscovering Continuations. Mathematical Structures in Computer Science, 9(4):367–401, 1999.

[San99b] D. Sangiorgi. The Name Discipline of Uniform Receptiveness. Theoretical Computer Science, 221(1–2):457–493, 1999. An abstract appeared in the Proceedings of ICALP ’97 , LNCS 1256, pages 303–313.

[San99c] D. Sangiorgi. Reasoning About Concurrent Systems Using Types. In W. Thomas, ed, Proceedings of FoSSaCS ’99, volume 1578 of LNCS, pages 31–40. Springer, 1999.

[San99d] D. Sangiorgi. The Typed π-Calculus at work: A Proof of Jones’s Parallelisation Theorem on Concurrent Objects. Theory and Practice of Object-Oriented Systems, 5(1), 1999. An early version was included in the Informal proceedings of FOOL 4, January 1997.

[San01a] D. Sangiorgi. Asynchronous process calculi: the first-order and higher-order paradigms (Tutorial). Theoretical Computer Science, 253(2):311–350, Feb. 2001.

[San01b] D. Sangiorgi. Extensionality and Intensionality of the Ambient Logics. In Proceedings of POPL ’01, pages 4–13. ACM, Jan. 2001.

[San02] D. Sangiorgi. Types, or: Where’s the Difference Between CCS and π? In L. Brim, P. Jančar, M. Křetinský and A. Kučera, eds, Proceedings of CONCUR 2002, volume 2421 of LNCS, pages 76–97. Springer, Aug. 2002.

[Sew97] P. Sewell. On Implementations and Semantics of a Concurrent Programming Language. In A. Mazurkiewicz and J. Winkowski, eds, Proceedings of CONCUR ’97, volume 1243 of LNCS, pages 391–405. Springer, 1997.

[Sew98] P. Sewell. Global/Local Subtyping and Capability Inference for a Distributed Picalculus. In K. G. Larsen, S. Skyum and G. Winskel, eds, Proceedings of ICALP ’98, volume 1443 of LNCS, pages 695–706. Springer, July 1998. Full version as Technical Report 435, Computer Laboratory, University of Cambridge.

[SK98] E. Sumii and N. Kobayashi. A Generalized Deadlock-Free Process Calculus. In U. Nestmann and B. C. Pierce, eds, Proceedings of HLCL ’98, volume 16.3 of ENTCS. Elsevier Science Publishers, 1998.

[SL97] J.-G. Schneider and M. Lumpe. Synchronizing Concurrent Objects in the Pi-Calculus. In R. Ducournau and S. Garlatti, eds, Proceedings of Langages et Modèles à Objects ’97, pages 61–76. Hermes, Roscoff, Oct. 1997.

[Smo94] G. Smolka. A Foundation for Higher-Order Concurrent Constraint Programming. In J.-P. Jouannaud, ed, Proceedings 1st International Conference of Constraints in Computational Logics, volume 845 of LNCS, pages 50–72. Springer, 1994. Available as Research Report RR-94-16 from DFKI Kaiserslautern.

[Sta96] I. Stark. A Fully Abstract Domain Model for the Pi-Calculus. In Proceedings of LICS ’96, pages 36–42. IEEE, Computer Society Press, July 1996.

[Ste94] G. Stefanescu. The Algebra of Flownomials Part 1: Binary Flownomials; Basic Theory. Report, Technical University Munich, Nov. 1994.

[Ste00] G. Stefanescu. Network algebra. Springer, 2000.

[SV99a] P. Sewell and J. Vitek. Secure Composition of Insecure Components. In Proceedings of the Computer Security Foundations Workshop, CSFW-12, 1999. Full version as Technical Report 463, Computer Laboratory, University of Cambridge, April 1999.

[SV99b] P. Sewell and J. Vitek. Secure Composition of Untrusted Code: Wrappers and Causality Types. Technical Report 478, Computer Laboratory, University of Cambridge, 1999.

[SV01] D. Sangiorgi and A. Valente. A distributed abstract machine for Safe Ambients. In Proceedings of ICALP 2001, volume 2076 of LNCS. Springer, July 2001.

[SW01] D. Sangiorgi and D. Walker. The π-calculus: a Theory of Mobile Processes. Cambridge University Press, 2001.

[SWP99] P. Sewell, P. Wojciechowski and B. Pierce. Location Independence for Mobile Agents. In H. Bal, B. Belkhouche and L. Cardelli, eds, Proceedings of ICCL ’98, Workshop on Internet Programming Languages (Chicago, IL, USA, May 13, 1998), volume 1686 of LNCS. Springer, Sept. 1999. Full version with title Location-Independent Communication for Mobile Agents: a Two-Level Architecture appeared as Technical Report 462, Computer Laboratory, University of Cambridge, April 1999.

[SY97] T. Sekiguchi and A. Yonezawa. A Calculus with Code Mobility. In Prodeedings of FMOODS’97, Canterbury, July. IFIP, Chapman and Hall, 1997.

[THK94] K. Takeuchi, K. Honda and M. Kubo. An interaction-based language and its typing system. In C. Halatsis, D. Maritsas, G. Philokyprou and S. Theodoridis, eds, Proceedings of PARLE ’94, volume 817 of LNCS, pages 398–413. Springer, 1994.

[Tho89] B. Thomsen. A Calculus of Higher Order Communicating Systems. In Proceedings of POPL ’89, pages 143–154. ACM, January 1989.

[Tho90] B. Thomsen. Calculi for Higher Order Communicating Systems. PhD thesis, Imperial College, University of London, Sept. 1990.

[Tho93] B. Thomsen. Plain CHOCS. A Second Generation Calculus for Higher Order Processes. Acta Informatica, 30(1):1–59, 1993.

[Tho95] B. Thomsen. A Theory of Higher Order Communicating Systems. Journal of Information and Computation, 116(1):38–57, 1995.

[TKLC95] B. Thomsen, F. Knabe, L. Leth and P.-Y. Chevalier. Mobile Agents Set to Work. Communications International, July 1995.

[TLK96] B. Thomsen, L. Leth and T.-M. Kuo. A Facile Tutorial. In U. Montanari and V. Sassone, eds, Proceedings of CONCUR ’96, volume 1119 of LNCS, pages 278–298. Springer, 1996.

[Tur96] D. N. Turner. The Polymorphic Pi-Calculus: Theory and Implementation. PhD thesis, LFCS, University of Edinburgh, June 1996. CST-126-96 (also published as ECS-LFCS-96-345).

[TZH02] D. Teller, P. Zimmer and D. Hirschkoff. Using Ambients to Control Resources. In L. Brim, P. Jančar, M. Křetinský and A. Kučera, eds, Proceedings of CONCUR 2002, volume 2421 of LNCS, pages 288–305. Springer, Aug. 2002.

[US01] A. Unyapoth and P. Sewell. Nomadic Pict: Correct Communication Infrastructures for Mobile Computation. In Proceedings of POPL ’01, pages 116–127. ACM, Jan. 2001.

[Var96] P. Varone. Implementation of “Generic Synchronization Policies” in Pict. Technical Report IAM-96-005, Software Composition Group, University of Berne, Apr. 1996.

[Vas92] V. T. Vasconcelos. Trace Semantics for Concurreny Objects. Master Thesis, Department of Computer Science, Keio University, Feb. 1992.

[Vas94a] V. T. Vasconcelos. Predicative Polymorphism in π-Calculus. In C. Halatsis, D. Maritsas, G. Philokyprou and S. Theodoridis, eds, Proceedings of PARLE ’94, volume 817 of LNCS, pages 425–437. Springer, 1994.

[Vas94b] V. T. Vasconcelos. A process-calculus approach to typed concurrent objects. PhD thesis, Keio University, 1994.

[Vas94c] V. T. Vasconcelos. Typed Concurrent Objects. In M. Tokoro and R. Pareschi, eds, Proceedings of ECOOP ’94, volume 821 of LNCS, pages 100–117. Springer, 1994.

[vB97] F. van Breugel. A Labelled Transition System for π_{ϵ}-Calculus. In M. Bidoit and M. Dauchet, eds, Proceedings of TAPSOFT ’97, volume 1214 of LNCS, pages 312– 336, Lille, France, Apr. 1997. Springer.

[VD98] J.-L. Vivas and M. Dam. From Higher-Order pi-Calculus to pi-Calculus in the Presence of Static Operators. In D. Sangiorgi and R. de Simone, eds, Proceedings of CONCUR ’98, volume 1466 of LNCS, pages 115–130. Springer, Sept. 1998.

[Ven97] A. Venet. Abstract Interpretation of the π-Calculus. In M. Dam, ed, Proceedings of the Fifth LOMAPS Workshop, volume 1192 of LNCS, pages 51–75. Springer, 1997.

[Ven98a] A. Venet. Automatic Determination of Communication Topologies in Mobile Systems. In G. Levi, ed, Proceedings of SAS ’98, volume 1503 of LNCS, pages 152–167. Springer, 1998.

[Ven98b] A. Venet. Automatic Determination of Communication Topologies in Mobile Systems. In G. Levi, ed, Proceedings of SAS ’98, volume 1503 of LNCS. Springer, 1998.

[VH93] V. T. Vasconcelos and K. Honda. Principal typing schemes in a polyadic π-calculus. In E. Best, ed, Proceedings of CONCUR ’93, volume 715 of LNCS, pages 524–538. Springer, 1993.

[Vic94] B. Victor. A Verification Tool for the Polyadic π-Calculus. Licentiate thesis, Department of Computer Systems, Uppsala University, Sweden, May 1994. Available as report DoCS 94/50.

[Vic98] B. Victor. The Fusion Calculus: Expressiveness and Symmetry in Mobile Processes. PhD thesis, Dept. of Computer Systems, Uppsala University, Sweden, June 1998.

[VLS98] V. T. Vasconcelos, L. Lopes and F. Silva. Distribution and Mobility with Lexical Scoping in Process Calculi. In U. Nestmann and B. C. Pierce, eds, Proceedings of HLCL ’98, volume 16.3 of ENTCS. Elsevier Science Publishers, 1998.

[VM94] B. Victor and F. Moller. The Mobility Workbench — A Tool for the π-Calculus. In D. Dill, ed, Proceedings of CAV ’94, volume 818 of LNCS, pages 428–440. Springer, 1994.

[VP96] B. Victor and J. Parrow. Constraints as Processes. In U. Montanari and V. Sassone, eds, Proceedings of CONCUR ’96, volume 1119 of LNCS, pages 389–405. Springer, 1996.

[VP98] B. Victor and J. Parrow. Concurrent Constraints in the Fusion Calculus. In K. G. Larsen, S. Skyum and G. Winskel, eds, Proceedings of ICALP ’98, volume 1443 of LNCS, pages 455–469. Springer, July 1998.

[VR99] V. T. Vasconcelos and A. Ravara. Comunication Errors in the π-Calculus are Undecidable. Information Processing Letters, 71:229–233, 1999.

[VT93] V. T. Vasconcelos and M. Tokoro. A Typing System for a Calculus of Objects. In S. Nishio and A. Yonezawa, eds, First JSSST International Symposion on Object Technologies for Advances Software (Kanazawa, Japan), volume 742 of LNCS, pages 460–474. Springer, 1993.

[Wal91a] D. Walker. π-calculus Semantics of Object-Oriented Programming Languages. In T. Ito and A. Meyer, eds, Proceedings of TACS ’91, volume 526 of LNCS, pages 532– 547. Springer, 1991. Available as Report ECS-LFCS-90-122, University of Edinburgh.

[Wal91b] D. Walker. Some Results on the π-Calculus. In A. Yonezawa and T. Ito, eds, Concurrency: Theory, Languages, and Architecture, volume 491 of LNCS, pages 21–35. Springer, 1991.

[Wal93] D. Walker. Process Calculus and Parallel Object-Oriented Programming Languages. In P. Tvrdik, T. Casavant and F. Plasil, eds, Parallel Computers: Theory and Practice, pages 369–390. Computer Society Press, 1993. Available as Research Report CS-RR-242, Department of Computer Science, University of Warwick.

[Wal94a] D. Walker. Algebraic Proofs of Properties of Objects. In D. Sannella, ed, Proceedings of ESOP ’94, volume 788 of LNCS, pages 501–516. Springer, 1994.

[Wal94b] D. Walker. On bisimulation in the π-calculus. In B. Jonsson and J. Parrow, eds, Proceedings of CONCUR ’94, volume 836 of LNCS, pages 315–330. Springer, Aug. 1994. Proceedings of 5th International Conference, Uppsala, Sweden.

[Wal95] D. Walker. Objects in the π-calculus. Journal of Information and Computation, 116(2):253–271, 1995.

[WB96] S. Weber and B. Bloom. Metatheory of the π-calculus. Technical Report TR96-1564, Cornell University, Jan. 1996.

[WG99] L. Wischik and P. Gardner. Symmetric Action Calculi (Abstract). In I. Castellani and B. Victor, eds, Proceedings of EXPRESS ’99, volume 27 of ENTCS. Elsevier Science Publishers, 1999.

[WS00] P. Wojciechowski and P. Sewell. Nomadic Pict: Language and Infrastructure Design for Mobile Agents. IEEE Concurrency, 8(2):42–52, 2000.

[YH99a] N. Yoshida and M. Hennessy. Assigning Types to Processes. CogSci Report 99.02, School of Cognitive and Computing Sciences, University of Sussex, UK, 1999.

[YH99b] N. Yoshida and M. Hennessy. Subtyping and Locality in Distributed Higher Order Processes. In J. C. Baeten and S. Mauw, eds, Proceedings of CONCUR ’99, volume 1664 of LNCS, pages 557–572. Springer, Aug. 1999.

[YHB02] N. Yoshida, K. Honda and M. Berger. Linearity and Bisimulation. In M. Nielsen and U. H. Engberg, eds, Proceedings of FoSSaCS 2002, volume 2303 of LNCS, pages 417–431. Springer, Apr. 2002.

[Yos95] N. Yoshida. Graph Notation for Concurrent Combinators. In T. Ito and A. Yonezawa, eds, Proceedings of TPPP ’94, volume 907 of LNCS, pages 393–412. Springer, 1995.

[Yos96] N. Yoshida. Graph Types for Monadic Mobile Processes. In V. Chandru and V. Vinay, eds, Proceedings of FSTTCS ’96, volume 1180 of LNCS, pages 371–386. Springer, 1996. Full version as Technical Report ECS-LFCS-96-350, University of Edinburgh.

[Yos98] N. Yoshida. Minimality and Separation Results on Asynchronous Mobile Processes: Representability Theorem by Concurrent Combinators. Technical Report 5/98, School of Cognitive and Computing Sciences, University of Sussex, UK, Sept. 1998. An extended abstract appeared in the Proceedings of CONCUR ’98, LNCS 1466.

[Zim00a] P. Zimmer. On the Expressiveness of Pure Mobile Ambients. In L. Aceto and B. Victor, eds, Proceedings of EXPRESS ’00, volume 39.1 of ENTCS, pages 81–104. Elsevier Science Publishers, 2000.

[Zim00b] P. Zimmer. Subtyping and Typing Algorithms for Mobile Ambients. In J. Tiuryn, ed, Proceedings of FoSSaCS 2000, volume 1784 of LNCS, pages 375–390. Springer, 2000.

[ZP97] X. Zhang and J. Potter. Class-based Models in the Pi-Calculus. In C. Mingins, R. Duke and B. Meyer, eds, Proceeding of The 25th International Conference in Technology of Object-Oriented Languages and Systems (TOOLS Pacific ’97, Melbourne, Australia), pages 219–231, November 1997.